### v1

# CytoNet: A Foundation Model for the Human Cerebral Cortex

[…]
CytoNet: A Foundation Model for the Human Cerebral Cortex
Christian Schiffer Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Zeynep Boztoprak Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Jan-Oliver Kropp Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Julia Thönnißen Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Katia Berr Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Hannah Spitzer Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Institute for Stroke and Dementia Research (ISD), LMU University Hospital, LMU Munich, Germany Katrin Amunts Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Timo Dickscheid Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Computer Vision, Institute for Computational Visualistics, University of Koblenz
Zusammenfassung
To study how the human brain works, we need to explore the organization of the cerebral cortex and its detailed cellular architecture. We introduce CytoNet, a foundation model that encodes high-resolution microscopic image patches of the cerebral cortex into highly expressive feature representations, enabling comprehensive brain analyses. CytoNet employs self-supervised learning using spatial proximity as a powerful training signal, without requiring manual labelling. The resulting features are anatomically sound and biologically relevant. They encode general aspects of cortical architecture and unique brain-specific traits. We demonstrate top-tier performance in tasks such as cortical area classification, cortical layer segmentation, cell morphology estimation, and unsupervised brain region mapping. As a foundation model, CytoNet offers a consistent framework for studying cortical microarchitecture, supporting analyses of its relationship with other structural and functional brain features, and paving the way for diverse neuroscientific investigations.
Recent advances in artificial intelligence have shown that large-scale foundation models can tackle problems once considered intractable, ranging from predicting protein structures [129] to powerful vision [131, 143] and language [139, 142] models. At the core of these successes is self-supervised learning, which extracts expressive features from massive unannotated datasets by generating implicit training signals from the data itself [122, 118]. This transition from narrow, task-specific systems to general-purpose representations marks a paradigm shift for computational problem-solving across science, technology, and industry.
[…]
Here we introduce CytoNet, a foundation model for human cortical organization trained with self-supervised learning on millions of microscopic image patches from over 4,000 histological sections of ten postmortem brains. CytoNet encodes local cytoarchitectonic patterns into expressive feature representations using the proposed SpatialNCE loss, a contrastive objective that leverages anatomical proximity: patches from nearby cortical locations are treated as similar, whereas those from distant locations are treated as dissimilar. This strategy captures biologically meaningful variation without manual annotations and yields a feature space that generalizes across brains and scales. CytoNet supports applications ranging from area and layer segmentation to unsupervised clustering and comparative analysis of cortical organization. By embedding image patches into a common feature space, it enables cross-brain comparisons that reveal both shared principles of cortical architecture and brain-specific differences.
1 Results
We present the results of CytoNet from three complementary perspectives. First, we present the pretraining strategy that allowed the model to learn from large-scale histological data without annotations. Second, we analyze the learned feature space, using dimensionality reduction and clustering metrics to assess how it reflects local, global, and inter-subject cytoarchitectonic variation. Third, we demonstrate its utility in downstream applications, including prediction of structural variation, cortical area and layer segmentation, and data-driven area discovery. Together, these results demonstrate CytoNet’s scalability and versatility, showing how a single representation supports diverse analyses.
1.1 CytoNet learns cortical cytoarchitecture from spatial proximity
Abbildung 1: Illustration of the self-supervised pretraining workflow using the proposed SpatialNCE loss in CytoNet. Spatial transformations between the MNI Colin 27 [83] 3D reference coordinate space (A) and microscopic scans of histological brain sections (C) of postmortem human brains (B) were used to link high-resolution microscopic image patches (E) with corresponding 3D locations in the common reference space, allowing to estimate distances between sampled image patches from different brains. These were used to compute similarity scores for the proposed SpatialNCE contrastive loss (D), which promotes extraction of expressive feature vectors for each image patch.
[…]
To train CytoNet, we introduced SpatialNCE, a contrastive objective that encourages patches from nearby cortical locations to map to similar representations. This leverages the anatomical continuity of the cortex as a heuristic for defining similarity, enabling scalable training on large unlabeled datasets without augmentations or manual annotations. SpatialNCE builds on InfoNCE [114], the basis of many modern contrastive methods [118, 122, 135, 117, 136], but uses distances in a shared coordinate space instead of image augmentations as proxy for semantic similarity.
In natural images, positive pairs are often created by augmentations that preserve semantic similarity while disrupting irrelevant features. For the type of data used for CytoNet, this assumption is problematic: common transformations can alter cytoarchitectonic structure, while confounding features such as blood vessels or folding geometry often remain unchanged [130, 124]. Supervised contrastive learning [123, 132] circumvents this issue by using labels, but is limited by annotation cost.
SpatialNCE overcomes both issues by defining similarity directly from anatomical proximity. Given a batch of image patches $x_{i}$ with corresponding 3D coordinates $p_{i}$ and normalized neural network features $z_{i}=f(x_{i})$, the loss for patch $i$ is:
$\mathcal{L}_{i}=-\frac{1}{\sum_{i\neq j}\omega_{ij}}\sum_{i\neq j}\omega_{ij}\log\frac{\exp\left(z_{i}^{\top}z_{j}/\tau\right)}{\sum_{k\neq i}\exp\left(z_{i}^{\top}z_{k}/\tau\right)},$ (1)
with similarity weights
[…]
All sections were co-registered to the MNI Colin 27 single subject reference space [83], enabling consistent distance computation across brains. Crucially, the model never receives spatial coordinates, and therefore similarity must be inferred from image content. Unlike augmentation-based methods that impose semantic invariance externally, SpatialNCE exploits the intrinsic continuity of brain organization, encouraging features that capture shared cytoarchitectonic properties (e.g., cell density, lamination) while suppressing confounds (e.g., staining variation or morphology).
The SpatialNCE loss can be used with any neural network architecture. We evaluated ResNet50 [106] and hybrid ResNet50-ViT-B [119] models with modified input layers to handle large input patches. To process the large datasets in reasonable timeframes (up to $4\text{\,}\mathrm{TB}\text{/}$ per epoch, $600\text{\,}\mathrm{TB}\text{/}$ per training), models were trained on 16 compute nodes (64 NVidia A100 $40\text{\,}\mathrm{GB}\text{/}$ GPUs) of the supercomputer JURECA-DC [134] at Jülich Supercomputing Centre, with a runtime of up to 28 hours.
1.2 CytoNet encodes cytoarchitectonic organization
Abbildung 2: Anatomical plausibility of feature representations learned by CytoNet-ViT (1M). Top: 2D UMAP plot of the learned latent space, color coded by maximum probability labels of corresponding coordinates in the Julich Brain atlas (version 3.1, [116]), as an approximate assignment to brain areas. Brain-specific clusters fan out along the first UMAP dimension, while the second UMAP dimension shows a transition from the occipital to the frontal pole. A gap along the anterior-posterior axis co-aligns with the central sulcus, marking a prominent structural and functional division. The cluster corresponding to B99 —not included during pretraining— appears more compact than the other clusters, but shows a comparable cytoarchitectonic organization. Bottom: Aggregated pairwise cosine similarity of features across ten brains. Cosine similarity was computed between feature vectors from image patches, grouped by Julich Brain Atlas labels and averaged over all area pairs. Rows and columns represent brain areas, ordered by hemisphere, lobe, and label; area names are omitted for clarity (see supplementary Table 3).
We examined how CytoNet-ViT (1M) organizes cytoarchitectonic information in its feature space (Figure 2, top [^1] Interactive versions of selected figures are available at https://go.fzj.de/cytonet-interactive.). CytoNet-ViT (1M) is a hybrid ResNet50–ViT-B model pretrained on one million cortical patches. 2D UMAP embeddings [109] of patches from all ten brains revealed distinct brain-specific manifolds (Calinski–Harabasz index ([77], CHI) 2517.84) with consistent internal organization: atlas labels indicating different cytoarchitectonic areas [116] clustered coherently (CHI 719.66). The second UMAP dimension showed a subdivision of clusters at the central sulcus, which is an important anatomical landmark separating motor and somatosensory areas. The tenth brain (B99), which was not included in the pretraining phase, appeared more compact than the remaining nine. Yet, it showed comparable internal structure, indicating generalization beyond the training set.
To assess how well CytoNet features express cytoarchitectonic similarity across brains, pairwise cosine similarities of feature vectors were mean-aggregated into similarity matrices grouped by atlas areas (Figure 2, bottom). Resulting matrices exhibited a strong block structure, i.e., there was higher similarity within areas than between them. These patterns were highly correlated across subjects (Pearson $r=0.88\pm 0.09$), confirming that CytoNet encodes stable inter-area relationships. The tenth brain again showed elevated overall similarity and weaker block structure, consistent with the reduced specificity observed in the UMAP embeddings.
Abbildung 3: Attention maps from the first self-attention layer of CytoNet-ViT (1M). The figure includes example patches from areas hOc1 (primary visual cortex, [88]), 4a (primary motor cortex, [82]), and 3b (primary somatosensory cortex, [86]). Each row shows the input image (left) and attention scores of all 12 heads overlaid on the image (red = stronger attention). Highlighted are the stripe of Gennari in the primary visual cortex (top), Betz giant cells in layer V of motor cortex (center), and a pronounced layer IV in somatosensory cortex (bottom). Attention scores were gamma transformed ($\gamma=0.5$) to aid visualization.
Further, we studied the attention maps of the class token in the first vision transformer layer to visualize how the model attends to cytoarchitectonic structures. Figure 3 shows attention scores for example patches from the primary visual cortex hOc1 (Brodmann area 17, [88]), the (anterior) primary motor cortex area 4a [82] and area 3b of the primary somatosensory cortex [86]. The attention scores reveal how the model attends to the composition of cortical layers that define cytoarchitecture. In particular, attention heads focusing on prominent landmarks like the stripe of Gennari in hOc1, Betzt giant cells in layer V of 4a, and a pronounced layer V in 3b were identified.
1.3 Applications of CytoNet
The representations learned by CytoNet provide a versatile foundation for multiple downstream applications in human brain mapping. To demonstrate their utility, we evaluated CytoNet features in three complementary settings: (1) correlation with established structural and cytoarchitectonic properties, (2) supervised mapping tasks such as brain area classification and cortical layer segmentation, and (3) exploratory analyses for data-driven discovery of new or refined cortical areas.
Predicting structural variation in cytoarchitecture
Abbildung 4: Comparison of predictive performance between intensity profiles and CytoNet-ViT (1M) features in B2020. A: Linear regression models using varying subsets of PCA components revealed substantially higher $R^{2}$ scores for CytoNet features compared to intensity profiles across all evaluated structural and morphological properties. Reported values reflect the average $R^{2}$ across 5-fold cross-validation. B: Absolute feature importance scores —derived from regression coefficients for the first 32 PCA components of CytoNet features— showed that components 1–3 were strongly associated with spatial location in MNI space and the density of cortical layer IV, while other properties were predominantly encoded in higher components. LI to L6 denote cortical layers I to VI. C: The cumulative explained variance across PCA components indicates that CytoNet features capture substantially more variance than intensity profiles.
We assessed which anatomical properties are reflected in the feature space of CytoNet-ViT (1M), focusing on morphological and cytoarchitectonic variables extracted from B2020, the BigBrain dataset [98]. The predictive power of CytoNet features was compared to that of intensity profiles extracted from the BigBrain dataset [112], which also aim to capture local cytoarchitectonic composition. Intensity profiles mainly reflected anterior–posterior position, whereas CytoNet captured all three spatial axes as well as morphological properties such as cortical thickness, curvature, and cutting angle (Figure 4, A). It also enabled accurate prediction of cortical layer thicknesses and layer-wise cell densities, which intensity profiles from the 20 micrometer model failed to represent consistently (Figure 4, A, left columns). Across all evaluated properties, higher-dimensional PCA projections of CytoNet features generally improved predictive accuracy, while performance of intensity profile projections remained largely constant (Figure 4, B). This indicates that CytoNet encodes a richer set of structural cues that are distributed across many dimensions of the feature space, enabling fine-grained modeling of complex cytoarchitectonic patterns. Consistent with this finding, the cumulative explained variance of the PCA projections was markedly higher for CytoNet features than for intensity profiles, suggesting that CytoNet learns a more structured and informative representation space for downstream analysis (Figure 4, C).
Supervised mapping of brain areas and cortical layers
We next evaluated whether CytoNet features support explicit mapping of cytoarchitectonic organization in supervised tasks. Two complementary settings were considered: (i) classification of cytoarchitectonic areas across multiple brains and (ii) segmentation of cortical layers within histological image patches.
Abbildung 5: Performance of cytoarchitectonic brain area classification using CytoNet. A: Macro-F1 scores obtained by linear probing of different models. Mean and standard deviation over three training runs are reported. See supplementary Table 5 for more detailed scores. If applicable, the number of pretraining samples are indicated after the model name. B: Distribution of prediction errors of CytoNet-ViT (1M) by error distance for seen, transfer, and unseen brains. Error distance was defined as the number of hops between the predicted and true brain area in the adjacency graph of the Julich Brain Atlas 3.1 [116], where 1-hop errors correspond to directly adjacent areas, and larger distances reflect increasing topological separation. C: Boxplots of the logit margins for CytoNet-ViT (1M) predictions stratified by error distance. The logit margin —the difference between the top two logits— serves as a proxy for model confidence and distance to the decision boundary [141]. Correct predictions reveal higher confidence, while incorrect predictions show decreasing confidence with increasing error distance.
In area classification, CytoNet consistently outperformed models trained from scratch, SimCLR [118], and supervised contrastive baselines [132] across seen, transfer, and unseen brains (Figure 5, A and supplementary Table 5, see supplementary Section 3.5 for model details). CytoNet-ViT (1M) reached the best scores on seen brains (macro-F1 0.69 with linear probing, 0.71 with finetuning), surpassing all supervised alternatives and SimCLR. On the transfer brain, where no area annotations were available for training, CytoNet again achieved the highest scores. Even on the unseen brain, which was excluded from both pretraining and classifier training, CytoNet models remained competitive, with compact CNN-based variants showing the strongest generalization. In contrast, SimCLR variants performed poorly across all splits, often below training from scratch. Retrieval analysis confirmed this shortcut learning, showing that SimCLR features clustered images by tissue morphology and vascular patterns rather than area identity (see supplementary Figure 10).
Error analysis (Figure 5, B) revealed that CytoNet’s misclassifications were largely confined to borders between adjacent areas: $\sim$80% of errors in seen and transfer brains occurred within 1 hop of the correct area in the atlas adjacency graph, and over 95% within 2 hops. For the unseen brain, these proportions were somewhat lower (52% and 85%), but errors remained topologically plausible. Confidence, measured by logit margin [141], was significantly higher for correct predictions and decreased systematically with error distance (Figure 5, F). These findings indicate that CytoNet’s errors mirror the uncertainties faced by human experts, with most mistakes arising at difficult borders rather than random misclassifications.
Abbildung 6: Results for cortical layer segmentation and data-driven discovery of brain areas. A: Macro-F1 scores from linear probing of different models with varying amounts of training data (mean ± SD across five-fold cross-validation; evaluated on a dedicated test set, see supplementary Table 7). If applicable, the number of pretraining samples are indicated after the model name. B: Example segmentation of cortical layers in area 4a (primary motor cortex, [82]) for scratch (200k) and CytoNet-ViT (1M) models for increasing fractions of the training set. C: K-means clusters (indicated by colors) of CytoNet-ViT (1M) features in the vicinity of frontal pole areas Fp1 and Fp2 in brain B66, with clusters pre-localized using Julich Brain Atlas 3.1 probability maps (threshold 50%) and annotations. Hemispheres are shown separately for visualization, but clustering was performed jointly.
Cortical layer segmentation further demonstrated the richness and data efficiency of CytoNet features. Using only 1% of the annotated dataset (7 training patches), linear probing of CytoNet-ViT (1M) reached a macro-F1 of 0.63, i.e., over four times the scratch baseline (0.15) and well above SimCLR-pretrained models (0.49). With 5% of training data, CytoNet already achieved macro-F1 of 0.74, while baselines required 20% of data to reach similar performance (Figure 6, A, B and supplementary Table 7). Finetuning CytoNet improved accuracy with larger training datasets but caused overfitting in small-data regimes and instability in transformer backbones, whereas linear probing remained robust in both cases.
Data-driven discovery of brain areas
We finally asked whether CytoNet features can support exploratory, data-driven refinement of brain parcellations through clustering. Accurately identifying and delineating brain areas remains a central challenge in brain mapping and analysis, particularly in regions lacking anatomical landmarks. While historical atlases such as Brodmann’s cytoarchitectonic map [73] provide a foundational parcellation, many regions have since been identified based on refined histological evidence, often by subdividing existing areas. As a case study, we examined the frontal pole, which was initially defined as Brodmann area 10, but later subdivided into Fp1 and Fp2 by [100]. We treated the region as if no subdivision were known and tested whether clustering in CytoNet’s feature space could recover the distinction.
CytoNet features from brain B66 were clustered using $k$-means, applied either to points pre-localized by probabilistic maps or to points annotated as Fp1/Fp2 (Figure 6, C). In the probabilistic setting, one cluster aligned well with Fp2, while Fp1 points split into multiple clusters with inter-hemispheric differences. This resulted in high purity with respect to anatomical labels ($0.97\pm 0.02$) and hemisphere ($0.91\pm 0.02$). In the annotation-based setting, clustering robustly separated Fp1 and Fp2, with mean purity $0.94\pm 0.001$. Here, hemisphere information was more evenly distributed across clusters (purity $0.52\pm 0.001$), indicating that areal identity dominated. Manually identifying cluster 0 with Fp1 and cluster 1 with Fp2 results (supplementary Figure 9) yielded an accuracy of $94.75\%$.
2 Discussion
We demonstrated that CytoNet learns cortical cytoarchitecture from large-scale histological data without requiring manual annotations. The model encodes laminar and areal organization into a feature space that generalizes across subjects, aligns with cortical morphology, and supports diverse applications including area classification, layer segmentation, and exploratory parcellation. These results show that anatomical proximity provides an effective training signal, turning spatial continuity into a self-supervised objective that captures the latent factors of cortical organization. Together, they establish CytoNet as a biologically grounded and practically scalable foundation for systematic analysis of cortical histology at unprecedented scale, capable of integrating cytoarchitectonic patterns across brains while preserving both global consistency and individual variability.
The properties of these representations can be traced back to the proposed SpatialNCE training strategy. SpatialNCE derives its training signal by assuming that spatially close samples are often semantically consistent. In terms of cortical organization, this translates to prior knowledge that textures observed in proximity often show architectural variations of the same functional modules. This principle distinguishes CytoNet from other self-supervised learning approaches, both contrastive [118, 123, 117, 122] and non-contrastive [121, 127]. It directly addresses the limitations we encountered with the SimCLR approach [118], which sometimes performed even below scratch training because it relied on augmentation-invariant confounding cues such as vascular patterns or local curvature rather than cytoarchitecture (shortcut learning, [120]). Why, then, does augmentation-based self-supervised learning like SimCLR succeed in natural images but fail here? As argued by [130], augmentations implicitly define which features are treated as task-relevant content and which as task-irrelevant style. In our setting, standard augmentations (e.g., intensity distortions, noise, geometric transforms) leave confounds like vascular patterns or folding geometry intact, encouraging the model to treat them as content. This limitation reflects both the dataset and the chosen patch size: While choosing sufficiently large fields of view to capture laminar profiles inevitably includes confounds, smaller inputs could remove confounds but sacrifice important context. On the other hand, stronger augmentations (e.g., elastic deformations) may lead to undesired distortions of cytoarchitectonic structure. From this point of view, proximity-based similarity measures serve as a biologically grounded augmentation, leveraging cortical continuity across 3D space and across subjects to generate natural variations that disrupt confounds while preserving cytoarchitecture. Beyond cytoarchitecture, SpatialNCE provides an intuitive and easy-to-use framework for multimodal self-supervision: any dataset anchored into a common spatial frame can be integrated under the same loss without modification, with only modest requirements for spatial alignment. Although demonstrated here for cortical cytoarchitecture, the principle generalizes to other domains where spatial proximity systematically relates to semantic structure (e.g., remote sensing), establishing it as a unifying paradigm for multimodal representation learning.
CytoNet’s training objective defines similarity based on the Euclidean distance between image patches. However, spatial coordinates are never provided as input to the model. Keeping this in mind, the organization of the learned feature space yields two key insights. First, the strong alignment between the learned feature space and anatomical location shows that local image information reliably reflects the spatial continuity of cortical organization. Second, the objective is at the same time flexible enough to deviate from this imposed geometry when needed: although its objective promotes similarity across brains, CytoNet recovers distinct brain-specific clusters with consistent internal organization. A striking example is the separation along the central sulcus: despite their adjacency in 3D space, the motor and somatosensory areas that are located anterior and posterior to the central sulcus form distinct regions in the feature space, illustrating how CytoNet recovers sharp cytoarchitectonic contrasts.
CytoNet identifies clusters with consistent internal structure within each brain, while at the same time revealing substantial differences between individuals. This pattern aligns with long-recognized interindividual variability in cytoarchitecture, first demonstrated in probabilistic maps of cytoarchitectonic areas [84] and conceptually framed as an essential feature rather than noise in [99]. Such variability can even exceed differences between areas [84], underscoring the need to capture subject-specific patterns in cortical organization. By providing brain-specific yet systematically comparable feature spaces, CytoNet offers a computational basis for analyzing this variability at scale and relating it to functional specialization, development, and disease.
This intriguing outcome raises a key question: how can a model trained with nothing more than a proximity-based contrastive loss recover precise and meaningful anatomical structure? A likely explanation lies in the neural network architecture: projection heads in contrastive learning act as information bottlenecks, suppressing augmentation-specific signals while preserving features aligned with the loss objective [118]. Consistent with this, our experiments (supplementary Section 7.2) showed that CytoNet’s projection head filters out brain-specific variation —still visible in backbone features— while preserving cytoarchitectonic structure that generalizes across subjects. However, projected features do not improve classification, indicating that the suppressed brain-specific variation is itself informative for distinguishing areas.
What exactly does CytoNet capture in its learned feature space, and how can these representations be understood? This question highlights the broader interpretability challenge that is common to deep learning methods. Classical profile-based features [76, 87], including those used here as baseline for predicting structural variations [112], are interpretable but also limited in scope. They mainly describe laminar intensity distributions while neglecting or oversimplifying cytoarchitectonic changes, such as cell columns, when moving across the cortical ribbon. CytoNet, in contrast, learns high-dimensional representations that encompass laminar profiles together with contextual factors such as curvature and thickness of the cortex. It also distinguishes effects of sectioning, for example different cutting angles that alter the apparent shape and size of layers and cells, and encodes approximate spatial location, which provides useful priors much like those used by human experts. The data-driven integration of these and other sources of variation in a coherent space makes it possible to disambiguate biological organization from technical influences and enables analyses that are more robust, transferable across datasets, and scalable to whole brains.
CytoNet provides a versatile foundation for automated brain mapping tasks, supporting applications such as brain area classification, cortical layer segmentation, and data-driven area discovery. Unlike earlier methods [111, 132] it generalizes across entire brains and multiple individuals without task-specific retraining, bringing fully automated mapping within reach. In practice, CytoNet outputs still require post-processing (e.g., spatial smoothing and topological constraints), but the model shifts the bottleneck from manual annotation to scalable, data-driven analysis, making it feasible to process entire brains, compare across individuals, and handle terabyte- to petabyte-scale datasets. To the best of our knowledge, CytoNet currently represents the most precise method for cytoarchitectonic area classification, particularly when applied to brains without annotated training data.
CytoNet not only outperformed existing computational methods for cytoarchitectonic area classification but also allows a valuable comparison with human expert mapping. Experts typically focus on selected areas and annotate sparsely, because their method relies on direct visual comparisons across adjacent regions to detect differences in layering and cellular composition. CytoNet, by contrast, analyzes isolated patches and scales to dozens of areas across complete brains. Both approaches, however, encounter their greatest challenges at areal borders, where cytoarchitectonic transitions follow complex and heterogeneous patterns that do not always align with sharp anatomical landmarks. Clustering experiments showed that these very transitions, while complicating border classification, can also reveal subdivisions, highlighting CytoNet’s potential for data-driven refinement. As datasets grow in resolution and size, such approaches will become increasingly important to complement expert-driven mapping. In future work, combining CytoNet with post-processing methods that enforce spatial smoothness and topological constraints [133] may further enhance its utility for atlas refinement.
Classification performance dropped when CytoNet was applied to brains excluded from pretraining (i.e., to the unseen brain). This suggests that models benefit from acquiring a representational fingerprint“ for each brain. What exactly constitutes this fingerprint —individual biases in cellular biology, unknown batch effects in histological processing, or other factors— remains an important question for future work. From a practical perspective, incorporating a new brain into pretraining is not difficult: It requires only to align digitized sections approximately to the common reference space, a procedure that is well understood and supported by image registration and anchoring tools. Most importantly, no manual annotations of the images are necessary. The practicality of including each new brain in the pretraining also depends on the pace at which brain scans are typically acquired, and on the computational requirements for training. In our lab, for example, new whole-brain histological datasets are typically acquired at a pace of one or two brains per year. CytoNet would thus need to be retrained at most two times per year (requiring approximately 3600 GPU-hours), which is practically feasible on a moderately sized GPU cluster.
Complementing brain area classification, CytoNet also achieved strong performance in cortical layer segmentation, even with as little as 7 annotated image patches for training. Despite the relatively simple setup of this study (i.e., one brain, minimal training), the predictive accuracy and data efficiency suggest that the benefits of pretraining extend beyond area classification. Based on these results, we believe that CytoNet provides an important first step towards scaling automated laminar mapping —pioneered in painstaking manual studies by Brodmann [73], von Economo [75], and the Vogts [74]— to the scale of whole brains and across individuals.
In summary, CytoNet represents a step towards a new generation of approaches to study brain organization that are: (i) anatomically rooted, capturing fundamental structural principles, (ii) scalable, enabling dense and reproducible mapping across whole brains, (iii) general, applicable across regions, subjects, and imaging modalities, and (iv) extensible, providing a foundation for multi-modal integration and holistic models of brain organization.
3 Methods
3.1 Microscopic images of histological human brain sections
The analyses were conducted using 4654 cell-body stained histological sections from ten postmortem human brains ($465\pm 19$ sections/brain, min 438, max 492) from the brain collections of our laboratories in Jülich and Düsseldorf, with a total dataset size of approximately $10.71$ terabytes. The processing protocol is detailed in [116] and briefly summarized in the following.
[…]
3.2 Generation of sampling locations in the cerebral cortex
A dataset for contrastive pretraining of CytoNet was prepared by defining sampling locations along the cortical midline in all brains. The cortical midline runs centered between the pial surface and the gray matter surface. Identifying the cortical midline based on 2D scans of histological sections is sometimes challenging due to the projection of the three-dimensional structure of the cortex onto the 2D image planes. To address this, each brain was approximately 3D reconstructed by section-to-section alignment, consistent midsurfaces in 3D were computed, and then projected back onto the 2D image planes  [^2] Code available at https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/brain3d..
[…]
These segmentation volumes were cleaned to remove segmentation errors, imprecise alignment, or histological artifacts. Tissue defects were detected by smoothing volumes with a median filter of size 3 in the posterior-anterior direction and computing the difference to the input volume. Larger tissue defects were identified by detecting large connected components in the difference volume, and then replaced by the result of the median filter. Small parts of detached tissue were removed by extracting all connected components that were smaller than 1% of the largest tissue component. Parts of the volume belonging to subcortical gray matter and the cerebellum, which are not handled by the employed segmentation pipeline, were manually identified and excluded using the 3DSlicer software [101]. The manual steps required approximately $30\text{\,}\min$ per brain.
[…]
The 2D midline through the cortex was derived by projecting the 3D midsurface back onto the 2D images. For each histological section, the plane that cuts through the midsurface at the location of the respective brain section was determined, and the intersection between this plane and the midsurface, which can be interpreted as virtually cutting“ the reconstructed brain, was computed. The intersection was transformed back onto the brain sections by inverting the transformations used for 3D reconstruction.
As a result of the smoothing and cleaning steps in 3D, points transformed from 3D to 2D were not always located exactly in the center of the cortex. To address this, a refinement step that pushes“ points towards the cortical midline was applied. For the refinement, the morphological skeleton of the cortex segmentations was derived. Laplacian fields between the skeleton and both the pial boundary and gray-white matter boundary were then computed using successive over-relaxation. Each point was then integrated through the gradient field of the Laplacian fields, limiting the maximum movement to $2\text{\,}\mathrm{mm}\text{/}$. Points that contained less than 50% tissue according to the tissue segmentation were excluded from further processing. In total, $4,546,775\text{\,}\mathrm{s}$ample points were created ($454,678\text{\,}\text{/}$\pm$44,339\text{\,}\text{/}$\$ points/brain, min $399,493\text{\,}\mathrm{,}$ max $539,030\text{\,}\mathrm{)}$.
During pretraining, the presented SpatialNCE loss requires each image patch to be associated with a corresponding spatial location in the brain for computing similarity between samples. To allow distance computation across brains, it is important for spatial locations to be defined in a common reference coordinate system. To accomplish this, we made use of the fact that digitized histological sections used for training CytoNet are a subset of the dataset that was used to create the Julich Brain Atlas [116], for which linear and non-linear transformations from the pixel space of the digitized histological sections and the individual brain template MNI Colin 27 [83] are available as part of the Julich Brain workflow. These transformations were used to associate each sampling location in the microscopic images with a corresponding location in the coordinate system of the MNI Colin 27 space. Details on the used transformation workflow are provided in [116]. In supplementary Section 7.4, we additionally evaluated models pretrained on coordinates from the MNI 152 ICBM 2009c Nonlinear Asymmetric template space [97], obtained by nonlinearly transforming coordinates from MNI Colin 27 space using siibra-python [144].
3.3 Deep neural network architectures
CytoNet was evaluted with two architecture variants: R50 and R50-ViT. R50 is a modified ResNet50 [106] architecture following [132], where the initial downsampling block (i.e., the first two convolutional layers and the pooling layer) is replaced with two convolutional layers ($5\times 5$ convolution with stride $4$ and $3\times 3$ convolution with stride 1, 64 filters each) and a $2\times 2$ maximum pooling operation to account for the significantly larger input image size compared to many other classification tasks. Each convolutional layer is followed by a batch normalization layer [104] and ReLU activation.
R50-ViT is a hybrid between R50 and the ViT-B vision transformer architecture [119], which is constructed by appending a ViT-B vision transformer to the feature map produced by R50. The transformer uses learned positional embeddings that are added to the incoming feature maps. A special class token is prepended to the transformer input sequence, which aggregates information from the entire input image. See [106] and [119] for an in-depth description of the ResNet50 and ViT-B neural network architectures, respectively.
Models trained using these architectures are referred to by a combination of the training paradigm, the model architecture, and the number of pretraining samples (if applicable). For brevity, the R50 part is omitted, as all evaluated models are either pure R50 architectures, or hybrids of R50 and ViT-B. For example, scratch refers to a model trained from scratch using the R50 architecture, SubCon-ViT refers to a model trained using supervised contrastive learning with the R50-ViT-B hybrid architecture, and CytoNet (1M) refers to CytoNet pretrained on 1 million samples using R50 architecture.
3.4 Self-supervised pretraining of CytoNet
Dataset
[…]
Training protocol
CytoNet was trained using Stochastic Gradient Descent (SGD) with Nesterov momentum and a momentum factor of 0.9 in combination with the LARS optimizer [108] with trust coefficient 0.02. The batch size was $B=2048$, and the learning rate was set to 0.01(B/256)=0.080.01*(B/256)=0.08, which was kept constant over the course of the training. Weight decay with a factor of 0.0001 was applied to all non-bias parameters of the model. The temperature parameter $\tau$ for contrastive pretraining was set to 0.07. Pretraining was performed for 150 epochs. The RBF kernel for the SpatialNCE loss used a bandwidth of $\sigma=$10\text{\,}\mathrm{mm}\text{/}$\$.
Data augmentation was applied to capture typical variations in the data, following the data augmentation strategy detailed in [132], which is briefly summarized below. Image patches were randomly rotated by $\theta\in U\left[-\pi,+\pi\right]$ ($U[a,b]$: uniform distribution over $[a,b]$), patch center positions were translated in a random direction by $d\sim U\left[0\textrm{mm},0.2\textrm{mm}\right]$, and mirrored vertically with a probability of $50\%$. Pixel intensities $x\in\left[0,1\right]$ were randomly augmented using unbiased gamma augmentation [107] $\alpha x^{\gamma}+\beta$ with parameters $\alpha\in U[0.9,1.0]$, $\beta\in U[-0.1,+0.1]$, $\gamma=\frac{\log\left(0.5+2^{-0.5Z}\right)}{\log\left(0.5-2^{-0.5Z}\right)}$, $Z\sim U[-0.05,+0.05]$. In addition, images were blurred with an isotropic Gaussian $G_{\sigma}(x)$ of kernel size $\sigma\sim U[0.125,1.0]$, or sharpened according to $x+\delta(G_{\sigma_{u}}(x)-x)$ with ${\sigma}_{u}\sim U[0.125,1.0]$, $\delta\sim U[0.5,1.5]$ with probability $25\%$, respectively.
Following [118], projection layers were attached to the respective backbone architecture and the contrastive loss was computed on the output of the projection layers. A fully-connected layer with as many hidden units as the respective backbone output (2048 for R50; 768 for R50-ViT), batch normalization [104], ReLU, and a final linear layer with $256$ units were applied to the output of global average pooling for R50 or the class token for R50-ViT.
3.5 Brain area classification
Dataset
The dataset used to train classifiers for 113 cytoarchitectonic areas was derived from annotations shown in the Julich Brain atlas (version 3.1, both hemispheres, [116]), following the protocol described in [132]. The list of areas is provided in supplementary Table 2. Annotations were available as contours outlining the outer boundaries of each area. To generate sampling locations for extracting image patches, the cortical midline was first computed from the morphological skeleton of the rasterized contours. Sampling points were then uniformly spaced along this midline at $1\text{\,}\mathrm{mm}\text{/}$ intervals.
From these potential sampling points, datasets for training, testing, and transferability evaluation were created. Brains were grouped into three categories —seen brains, transfer brains, and unseen brains— based on their inclusion in (i) self-supervised pretraining (if applicable) and (ii) supervised training for brain area classification. Our default configuration for these datasets is shown in Table 1. Brains were randomly assigned to one of the three categories.
Sections from seen brains were divided into training and test sections, with 80% used for supervised training and 20% reserved for testing. Test sections were selected by choosing every fifth annotated section from each brain. This setup reflects a realistic use case in which models are applied to new sections from a brain with existing partial annotations.
Transfer brains were included in self-supervised pretraining but excluded from supervised training. This configuration enables evaluation of how well the learned representations transfer to brains without annotated training data, a common scenario in ongoing brain mapping efforts where new brains are digitized but not yet manually labeled.
Unseen brains were excluded from both self-supervised and supervised training. This strict separation provides a measure of generalization to entirely novel brains.
To address class imbalance in the training set —caused by variations in the size of cytoarchitectonic areas— stratified sampling with replacement was applied. For each area, 1200 image patches were sampled, resulting in a balanced training set of $135,600\text{\,}\mathrm{p}$atches. The sampling rate was chosen such that the median ratio between sampled and available patches across all areas was close to 1. Test datasets were not resampled and reflect the natural distribution of area sizes.
Tabelle 1: Brains used for pretraining, linear probing and finetuning.
dataset brain(s) pretraining supervised training
seen brains B11, B33, B44, B55
B66, B1010, B1212, B2020
transfer brain B77
unseen brain B99
Training protocol
Models were trained to classify 113 cytoarchitectonic brain areas by attaching a linear classifier to R50 or R50-ViT architectures (Section 3.3). Projection layers from pretrained models were discarded and replaced with a single linear layer (without bias) consisting of 113 output units. This classifier was attached to the global average pooling output (for R50) or the class token (for R50-ViT). Several baseline models were compared to CytoNet:
  • Training from scratch: full supervised training from randomly initialized weights.
  • Supervised contrastive learning (SupCon) [123, 132]: supervised contrastive pretraining on labeled training samples.
  • SimCLR (200k and 1M) [127]: self-supervised pretraining from semantic consistency under multi-view augmentation.
  • CytoNet (200k and 1M): self-supervised pretraining from spatial consistency in 3D space using SpatialNCE loss.
Classifier training for pretrained models (SupCon, SimCLR, CytoNet) was performed for 30 epochs, while training from scratch was performed for 180 epochs. Note that the number of epochs is not directly comparable across all models, since training from scratch and SupCon pretraining are limited to annotated samples, while SimCLR and CytoNet make use of unannotated samples as well. Nevertheless, the chosen number of epochs was sufficient to ensure convergence in both pretraining and supervised training.
For all models, two training strategies were evaluated:
  • Linear probing: Only the classifier was trained, while the pretrained backbone and corresponding batch normalization statistics remained frozen. Note that this setting is not applicable for training from scratch.
  • Finetuning: Both the classifier and backbone weights were optimized jointly.
All models were trained using categorical cross-entropy loss and the same data augmentation protocol as used during CytoNet pretraining (Section 3.4).
Unless otherwise noted, models were trained using SGD with Nesterov momentum and a scaled learning rate of $0.01\times(B/256)=0.08$, with batch size $B=2048$, and weight decay of 0.0001 applied to all non-bias parameters. Model-specific adjustments were necessary to stabilize training and included:
  • scratch-ViT and CytoNet-ViT (200k): learning rate reduced to 0.008.
  • SupCon-ViT: trained with AdamW and learning rate 0.001.
To evaluate the effect of the projection layer, we also trained and tested a variant of CytoNet-ViT (1M), denoted as CytoNet-ViT (1M) P, where the classifier was attached to the output of the pretrained projection layer. Both linear probing and finetuning were performed to assess whether the projected feature space contains sufficient task-relevant information.
3.6 Cortical layer segmentation
Dataset
[…]
The dataset was split into a fixed 80-20 split (732 training and 184 test patches). To simulate varying annotation budgets, 1%, 5%, 10%, 20%, and 100% subsets of the training pool were sampled. Each training subset was further split into five folds. For each fold, a model was trained and then evaluated on the fixed test set. This repeated sampling design allowed estimating the stability of model performance across different training subsets of equal size. All results are reported on the held-out test set. Mean and standard deviation of class-wise macro-F1 scores across the five trained models for each configuration are reported. To ensure a fair evaluation, mirror padding used to make irregularly shaped images compatible with the models was removed before computing scores.
Training protocol
Models were trained to classify the layer structure of the isocortex by attaching a pixel-wise linear classifier to R50 or R50-ViT architectures (Section 3.3). Projection layers from pretrained models were discarded and replaced with a $1\times 1$ convolutional layer (without bias) with 7 channels (six layers plus background). This classifier was attached to the global average pooling output (for R50) or the class token (for R50-ViT). Similar to brain area classification (Section 3.5), training from scratch, SimCLR (200k and 1M) [127], and CytoNet (200k and 1M) were compared, each trained using linear probing and finetuning. All models used either R50 or R50-ViT architectures, AdmaW optimizer with a learning rate of 0.001 and weight decay of 0.01, and categorical cross-entropy loss. No data augmentation was applied during training.
3.7 Predicting structural variations in cytoarchitecture
Extraction of morphological features from BigBrain
Morphological features for brain B2020 were extracted using the 3D reconstruction of the BigBrain dataset [116]. Cortical curvature was computed using the highres-cortex [105] module of the BrainVisa software [96] based on the gray-white matter segmentation available through the siibra tool suite [144]. To assess the cutting direction, the angle between the histological cutting plane and the cortical surface normal was measured. Surface normals were approximated by computing the gradient of a Laplace field defined between the pial surface and the gray–white matter boundary. The cutting direction was then quantified as the angle between these gradient vectors and the anterior–posterior axis, which corresponds to the histological cutting direction in BigBrain. Low angles indicate near-orthogonal slicing relative to the cortical sheet, whereas high angles reflect more oblique cuts, which can obscure the cortical lamination pattern [87]. Cortical thickness and layer-specific thicknesses (layers I–VI) were obtained from surface meshes described in [125], accessible via the siibra suite. Thickness was computed as the Euclidean distance between corresponding mesh vertices across laminar surfaces and mapped to the closest points on the cortical midline. Finally, computed features were assigned to points sampled along the cerebral cortex (Section 3.2) based on location in BigBrain space. All computed features were assigned to sampled cortical points (Section 3.2) based on their coordinates in BigBrain space.
Extraction of layer-wise cell-density from microscopic image patches
[…]
BigBrain intensity profiles
Intensity profiles [138] sampled across the BigBrain dataset [98] capture depth-dependent structural variations by measuring pixel intensity gradients across the cortical sheet. A publicly available dataset from [138] was used, comprising 327684 cortical profiles. Each profile contains 200 equidistant intensity values sampled along a 1-pixel wide line perpendicular to the cortical surface, extending from the pial boundary to the gray–white matter interface. Prior to correlation analysis, all intensity values were standardized via z-scoring.
Correlative analysis of CytoNet features and Bigbrain intensity profiles
To assess the information encoded in CytoNet representations, we conducted a correlative analysis between feature embeddings and known morphological and structural properties of the cerebral cortex. Learned feature representations of CytoNet-ViT (1M) were compared with structural features extracted from brain B2020, including cortical curvature, cortical thickness, layer-specific thickness, cutting direction, and layer-wise cell densities. In parallel, we conducted the same analysis using depth-wise intensity profiles (see above) to serve as a baseline for comparison.
For both CytoNet features and intensity profiles, correlations against all target features were evaluated using two input feature sets: the raw features, and principal component projections with dimensionalities 1, 16, 32, 64, 128, and 256. Principal component analysis (PCA) was fitted once on the entire dataset and applied consistently across all folds. To quantify predictive strength, separate linear regression models for each target feature were trained. Models were evaluated using five-fold cross-validation, with $R^{2}$ scores computed on the held-out test folds and reported as mean $\pm$ standard deviation across folds. All regressions were implemented using scikit-learn with default hyperparameters. $R^{2}$ scores for the prediction of layer-wise cell density using CytoNet features are not available, as the number of samples in each fold (730) was not sufficient to fit a linear classifier based on all dimensions of the CytoNet features (768).
To further interpret the latent structure of CytoNet representations, feature attribution was performed for the first 16 principal components of the CytoNet feature space. Linear attribution weights were computed by fitting a separate linear regressor to each target property and normalizing the regression coefficients by the ratio of input to output standard deviation. This normalization ensures that attributions are comparable across targets with differing dynamic ranges.
To evaluate the generalizability of spatial encoding in CytoNet representations, regressors were trained on data from brain B2020 and applied to spatial locations in other brains (supplementary Table 4). No finetuning or domain adaptation was performed across brains in this analysis.
3.8 Data-driven parcellation of cortical areas
Dataset
We assessed whether CytoNet feature embeddings support data-driven subdivision of cortical areas by focusing on the frontal pole regions Fp1 and Fp2 [100] in B66. These two areas form subdivisions of Brodmann area 10 and are located in the anterior portion of the prefrontal cortex. To pre-localize the region of interest, two complementary sources of anatomical information with different levels of spatial precision were used: (1) image-level annotations of Fp1 and Fp2 from [100], and (2) probabilistic maps for the same areas provided by the Julich Brain Atlas 3.1 [116]. Cortical sampling points were generated along the midline as described in Section 3.2, without spatial subsampling. Each point was assigned both a discrete label based on the image-level annotations and a probabilistic value from the atlas maps.
Clustering
$k$-means clustering was used to evaluate whether CytoNet feature representations can differentiate between areas Fp1 and Fp2 of the frontal pole [100]. CytoNet features were extracted from the backbone of the trained R50-ViT (1M) model, prior to the final projection head, and used directly without PCA or normalization. Clustering was performed separately for two subsets of cortical points, corresponding to the two pre-localization strategies.
For the atlas-based approach, all points with a probability greater than 50% of belonging to either Fp1 or Fp2 were selected. These points were clustered into $k=6$ groups to account for the spatial uncertainty and potential heterogeneity of the probabilistic maps.
For the annotation-based approach, only points that were explicitly labeled as Fp1 or Fp2 were selected. These were clustered into $k=2$ groups to match the number of known areas and to simulate the task of deciding whether a previously identified region should be subdivided.
In both settings, clustering was performed across points from both hemispheres using Euclidean distance in the CytoNet feature space. No feature normalization or spatial smoothing was applied prior to clustering. Cluster identity was not constrained by anatomical proximity or continuity. The consistency of cluster-to-label alignment was quantified using cluster purity. The purity of a cluster is defined as the fraction of its samples belonging to the most frequent ground truth label. For each clustering run, the mean purity across all clusters was computed. The overall clustering performance is reported as the mean and standard deviation of these average purities across 30 runs.
3.9 Computational setup
CytoNet pretraining (Section 3.4) and linear probing for brain area classification (Section 3.5) were performed on 16 compute nodes of the supercomputer JURECA-DC [134] at Jülich Supercomputing Centre (JSC, Forschungszentrum Jülich, Jülich, Germany). Each compute node was equipped with four Nvidia A100 GPUs ($4\times 6912$ CUDA cores, $4\times 432$ tensor cores, $4\times 40GB$ HBM2e memory), two AMD EPIC 7742, $2\times 64$ cores à 2.5GHz with hyperthreading, 512 GB memory, and InfiniBand HDR100 interconnect. Training, inference, and evaluation were implemented using ATLaS [^4] https://jugit.fz-juelich.de/inm-1/bda/software/analysis/atlas/atlas, a Python framework that enables large-scale neural network training for high-resolution microscopic image data.
ATLaS uses PyTorch [115] for neural network training. Training was parallelized using distributed data parallel (DDP) training, where each GPU processes a subset of samples in each batch and averages gradients before applying parameter updates. During contrastive pretraining, features computed by each GPU were gathered to compute pairwise similarities across all samples of a batch. Statistics computed by batch normalization layers were synchronized across GPUs in each training step. Automatic Mixed Precision (AMP) was applied to improve training performance by computing certain compute operations with reduced floating point precision. Gradient checkpointing was used for training R50-ViT models, setting checkpoints after all residual blocks of the convolutional network except for the first one, as well as all transformer layers. The average GPU memory footprint during contrastive pretraining (32 samples per GPU) was 37.6 GiB for R50 and 34.4 GiB for R50-ViT (with gradient checkpointing). Pretraining (Section 3.4) on 1 million samples for 150 epochs took approximately 28h (1792 GPU hours, 75 GPU days). Linear probing or finetuning for brain area classification (Section 3.5) on 131250 samples for 30 epochs took approximately 1.5 hours (96 GPU hours, 4 GPU days).
Training of models for cortical layer segmentation was performed on a workstation equipped with an Nvidia RTX 4090 GPU (24 GB RAM), Intel i9-14900k (32 cores à 6 GHz), and 192 GB RAM, taking between 5 minutes and 40 minutes per fold, depending on the fraction of used training samples, the model architecture, and whether linear probing or finetuning was used. The linear regression of structural variations (Section 3.7) was performed on the same workstation, taking up to 30 seconds per fold, depending on the dimensionality of the input vectors. No GPU acceleration was used for linear regression.
3.10 Use of Large-Language Models
ChatGPT (version 5) was used to prepare the manuscript to improve the clarity and brevity of the text. All generated text was revised, checked for correctness, and accepted by all authors.
4 Acknowledgement
This project received funding from the European Union’s Horizon 2020 Research and Innovation Programme, grant agreement 101147319 (EBRAINS 2.0 Project), the Helmholtz Association port-folio theme “Supercomputing and Modeling for the Human Brain”, the Helmholtz Association’s Initiative and Networking Fund through the Helmholtz International BigBrain Analytics and Learning Laboratory (HIBALL) under the Helmholtz International Lab grant agreement InterLabs-0015, from HELMHOLTZ IMAGING, a platform of the Helmholtz Information & Data Science Incubator [X-BRAIN, grant number: ZT-I-PF-4-061], and from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 46/1 – 501864659. Computing time was granted through JARA on the supercomputer JURECA-DC at Jülich Supercomputing Centre (JSC). The authors declare no competing interests.
5 Ethics declaration
The presented study requires no separate ethical approvals. All usage in this work is covered by a vote of the ethics committee of the Medical Faculty of the Heinrich Heine University Düsseldorf (#4863). Postmortem brains were obtained through body donor programs of the anatomical institutes of the universities of Düsseldorf, Rostock, and Aachen, in accordance with legal and ethical regulations and guidelines. All body donors have signed a declaration of agreement.
6 Data and code availability
All source code is publicly available under Apache 2.0 licence. The source code used for model training and evaluation is at https://jugit.fz-juelich.de/inm-1/bda/software/analysis/atlas/atlas. A docker container for reproducible execution is provided at https://jugit.fz-juelich.de/inm-1/bda/software/analysis/atlas/atlas_container. Trained model weights, corresponding configuration files, and feature embeddings are available at https://jugit.fz-juelich.de/inm-1/bda/software/analysis/cytonet_model_zoo. Code for approximate 3D reconstructions is available at https://jugit.fz-juelich.de/inm-1/bda/software/data_processing/brain3d. Code for tissue segmentation is available at https://jugit.fz-juelich.de/inm-1/bda/software/analysis/tseg. Due to ongoing processing, the full histological image dataset is not yet publicly available. However, selected sections are accessible via the EBRAINS Knowledge Graph (https://doi.org/10.25493/JWTF-PAB). Interactive versions of selected figures are available at https://go.fzj.de/cytonet-interactive.
7 Supplementary material
7.1 List of cytoarchitectonic areas
Tabelle 2: List of 113 brain areas used for brain area classification. Areas are denoted by the nomenclature of the Julich Brain Atlas [116], e.g., hOc1 for human occipital area 1 or FG1 for fusiform gyrus area 1.
occipital lobe
hOc1 hOc2 hOc3v hOc4v hOc3d hOc4d hOc4la hOc4lp
hOc5 hOc6
parietal lobe
Ip1 Ip2 Ip3 Ip4 1 2 3a 3b
5L 5M 5Ci 7PC 7A hIP3 PF PFcm
PFm PFop PFFt PGa PGp hIP1 hIP2 hIP4
hIP5 hIP6 hIP7 hIP8 hPO1
temporal lobe
FG1 FG2 FG3 FG4 Te 1.0 Te 1.1 Te 1.2 Te 2.1
Te 2.2 Te 3 STS1 STS2 TeI TI
insula
Ig1 Ig2 Id1 Ig3 Id2 Id3 Id4 Id5
Id6 Ia7 Ia1
frontal lobe
4a 4p 6d1 6d2 6d3 6v1 6v2 6r1
6mp 6ma 11a 11p 13 Fo4 Fo5 Fo6
Fo7 IFJ1 IFJ2 IFS1 IFS2 IFS3 IFS4 8a
8b 8c 8d SFS1 SFS2 FMS1 MFG1 44
45 Op5 Op6 Op7 Op8 Op9
limbic lobe
25a 25p s24a s24b s32 p24a p24b pv24c
pd24cd pd24cv p32
Tabelle 3: List of areas from the Julich Brain Atlas (version 3.1, [116]) used in Figure 2. For brevity, pre- and postfixes to the area name are omitted (e.g., Area hOc1 (V1, 17, CalcS)“ is abbreviated as hOc1“). Indices of areas in the similarity matrices in Figure 2 are provided.
occipital lobe
0 hOc1 3 hOc3v 6 hOc4lp 9 hOc6
1 hOc2 4 hOc4d 7 hOc4v
2 hOc3d 5 hOc4la 8 hOc5
parietal lobe
10 1 18 7M 26 PFcm 34 hIP3
11 2 19 7P 27 PFm 35 hIP4
12 3a 20 7PC 28 PFop 36 hIP5
13 3b 21 Op1 29 PFt 37 hIP6
14 5Ci 22 Op2 30 PGa 38 hIP7
15 5L 23 Op3 31 PGp 39 hIP8
16 5M 24 Op4 32 hIP1 40 hPO1
17 7A 25 PF 33 hIP2
temporal lobe
41 CoS1 47 OTS1 53 TI 59 Te 2.2
42 FG1 48 Ph1 54 TPJ 60 Te 3
43 FG2 49 Ph2 55 Te 1.0 61 TeI
44 FG3 50 Ph3 56 Te 1.1
45 FG4 51 STS1 57 Te 1.2
46 FG5 52 STS2 58 Te 2.1
insula
62 Ia1 66 Id10 70 Id5 74 Id9
63 Ia2 67 Id2 71 Id6 75 Ig1
64 Ia3 68 Id3 72 Id7 76 Ig2
65 Id1 69 Id4 73 Id8 77 Ig3
frontal lobe
78 44 90 6v3 102 Fp1 114 Op10
79 45 91 8d1 103 Fp2 115 Op5
80 4a 92 8d2 104 IFJ1 116 Op6
81 4p 93 8v1 105 IFJ2 117 Op7
82 6d1 94 8v2 106 IFS1 118 Op8
83 6d2 95 Fo1 107 IFS2 119 Op9
84 6d3 96 Fo2 108 IFS3 120 SFG2
85 6ma 97 Fo3 109 IFS4 121 SFG3
86 6mp 98 Fo4 110 MFG1 122 SFG4
87 6r1 99 Fo5 111 MFG2 123 SFS1
88 6v1 100 Fo6 112 MFG4 124 SFS2
89 6v2 101 Fo7 113 MFG5
limbic lobe
125 33 128 s32 131 CA3 134 TrS
126 EC 129 CA1 132 DG 135 Tu
127 p32 130 CA2 133 HATA 136 TuTi
Table 3 shows the list of areas from the Julich Brain Atlas (version 3.1, [116]) used for analyses presented in Figure 2. Table 2 provides the names for the 113 areas used in cytoarchitectonic classification, selected as the subset of atlas annotations available in at least four of the ten brains analyzed in this study.
7.2 Extended embedding analysis
Abbildung 7: 2D UMAP embeddings from features before (left) and after (right) the projection layers of CytoNet-ViT (1M). Embeddings are color coded by brain (A, B), Julich Brain labels (C, D), and hemisphere (E, F). Features extracted before the projection layers form brain-specific clusters in the UMAP space (A) with similar internal arrangement of atlas labels (C) and hemispheres (E). In comparison, features extracted after the projection layers show strong alignment between points from different brains (B) with consistent atlas labels (D) and hemispheres (F). The transfer brain B99 appears more compact and less differentiated than other brains, both before (A) and after the projection (B).
Following common practice in self-supervised learning [118], CytoNet was trained with a shallow projection head appended to the backbone. We compared the latent space structure of features extracted from the backbone (our default), and after the shallow projection head of CytoNet-ViT (1M). Backbone features preserved brain identity (Figure 7A,C,E), whereas projected features showed minimal brain separation and instead aligned strongly with hemisphere (CHI 11577.88 → 28416.23) and atlas label (CHI 709.04 → 3648.73). Thus, the projection layers effectively suppressed brain-specific artifacts (e.g., staining, morphology, sectioning) and emphasized spatially consistent cytoarchitectonic organization. In the held-out brain, projections retained a global anterior–posterior structure but showed weaker differentiation by atlas label (CHI $77.63$ vs. $122.96\pm 12.17$) and hemisphere (CHI $27.85$ vs. $1633.94\pm 231.41$), reflecting reduced feature specificity.
Abbildung 8: 3D UMAP plots of CytoNet-ViT (1M) features, colored by hemisphere (A), posterior-anterior location (B) and inferior-superior location (C) in the MNI Colin 27 space. Embedding dimensions strongly correlated with spatial locations, forming distinct hemisphere-specific manifolds.
Complementing our 2D UMAP analyses (Figure 2), we visualized 3D UMAP embeddings colored by hemisphere, anterior-posterior and inferior-superior axes (Figure 8). The results show strong alignment between embedding locations and spatial locations, including distinct hemisphere-specific manifolds.
7.3 Cross-brain prediction of spatial coordinates
To assess how well CytoNet features generalize across brains in terms of spatial encoding, linear regression models were trained to predict MNI coordinates from feature representations. Specifically, one model was fitted per spatial axis (anterior-posterior, inferior-superior, and left-right) using features extracted from brain B2020. Models were trained using 5-fold cross-validation on B2020, and then applied to predict spatial coordinates in other brains based on their CytoNet-ViT-B (1M) features. Euclidean prediction errors were computed in MNI Colin 27 reference space (Table 4), where the postmortem brains are being presented after registration [116]. This approach allows evaluating how consistently CytoNet encodes spatial location across different individuals. Prediction errors were approximately $10\text{\,}\mathrm{mm}\text{/}$ in anterior-posterior and inferior-superior directions. Errors in left-right direction were larger (approximately $20\text{\,}\mathrm{mm}\text{/}$), likely due to structural left-right symmetries. Substantially higher errors were observed in the unseen brain B99, indicating reduced spatial generalization without brain-specific pretraining. Such cross-brain prediction experiments can serve as a proxy for evaluating the general utility of learned features for other structural properties, such as cortical thickness, layer boundaries, or cell density, which are often spatially organized and may benefit from similarly aligned representations.
Tabelle 4: Euclidean distance between true and predicted locations in MNI Colin 27 space. Linear models were fitted on CytoNet-ViT (1M) features from B2020 and applied to features from other brains. Mean and standard deviation across 5-fold cross validation on B2020 are shown. Prediction errors in anterior-posterior and inferior-superior directions were approximately $10\text{\,}\mathrm{mm}\text{/}$, and approximately $20\text{\,}\mathrm{mm}\text{/}$ in left-right direction. Prediction errors in the unseen brain B99 were significantly higher than brains that were included for pretraining of CytoNet-ViT (1M).
dimension left-right anterior-posterior inferior-superior
B01 $9.62\pm 0.12 11$$1.56\pm 0.11$ $9.89\pm 0.12$
B03 $9.08\pm 0.08 11$$1.97\pm 0.15$ $10.92\pm 0.10$
B04 $18.60\pm 0.09$ $12.19\pm 0.12$ $9.98\pm 0.04$
B05 $20.10\pm 0.29$ $12.87\pm 0.08$ $9.79\pm 0.09$
B06 $18.01\pm 0.43$ $12.97\pm 0.06$ $9.79\pm 0.05$
B07 $21.16\pm 0.13$ $15.91\pm 0.18$ $12.04\pm 0.26$
B09 $42.67\pm 0.35$ $22.42\pm 0.08$ $16.19\pm 0.11$
B10 $20.09\pm 0.16$ $12.93\pm 0.17$ $9.73\pm 0.03$
B12 $19.08\pm 0.26$ $10.80\pm 0.04$ $9.65\pm 0.10$
7.4 Extended scores for brain area classification and layer segmentation
Tabelle 5: Scores for brain area classification obtained by different models using linear evaluation and finetuning. Mean and standard deviation of scores across three training runs with different random initialization are reported. The model suffixed with P was trained on outputs of the projection layer used during contrastive learning rather than backbone features. The model suffixed with M was pretrained using spatial coordinates from the ICBM 152 space rather than MNI Colin 27 reference template.
linear probing (encoder frozen) finetuning (encoder trainable)
seen brains macro-F1 top-1 acc top-3 acc macro-F1 top-1 acc top-3 acc
scratch - - - $0.39\pm 0.33$ $0.45\pm 0.32$ $0.64\pm 0.43$
scratch-ViT - - - $0.60\pm 0.01$ $0.65\pm 0.01$ $0.87\pm 0.01$
SimCLR (200k) $0.24\pm 0.00$ $0.33\pm 0.00$ $0.55\pm 0.00$ $0.33\pm 0.28$ $0.36\pm 0.31$ $0.55\pm 0.46$
SimCLR (1M) $0.15\pm 0.01$ $0.25\pm 0.01$ $3\pm 0.01 0.44±0$$4\pm 0.01$ $0.51\pm 0.01$ $0.78\pm 0.02$
SimCLR-ViT (200k) $0.06\pm 0.00$ $0.17\pm 0.01$ $0\pm 0.01 0.33±0$$3\pm 0.09$ $0.39\pm 0.09$ $0.67\pm 0.10$
SimCLR-ViT (1M) $0.05\pm 0.00$ $0.16\pm 0.01$ $0.29\pm 0.01$ $0.37\pm 0.06$ $0.45\pm 0.06$ $0.72\pm 0.06$
SupCon $1\pm 0.00 0.66±0$$6\pm 0.00$ $0.91\pm 0.00$ $0.59\pm 0.01$ $0.65\pm 0.01$ $0.89\pm 0.01$
SupCon-ViT $0.60\pm 0.00$ $0.65\pm 0.00$ $0.91\pm 0.00$ $0.51\pm 0.08$ $0.57\pm 0.08$ $0.84\pm 0.07$
CytoNet (200k) $0.64\pm 0.00$ $0.69\pm 0.00$ $0.93\pm 0.00$ $0.64\pm 0.01$ $0.69\pm 0.01$ $0.92\pm 0.01$
CytoNet (1M) $0.54\pm 0.00$ $0.62\pm 0.00$ $0.90\pm 0.00$ $0.67\pm 0.02$ $0.72\pm 0.01$ $0.94\pm 0.01$
CytoNet-ViT (200k) $0.54\pm 0.01$ $0.61\pm 0.01$ $0.90\pm 0.01$ $0.67\pm 0.00$ $0.72\pm 0.00$ $0.94\pm 0.00$
CytoNet-ViT (1M) $\mathbf{0.69}\pm 0.00$ $\mathbf{0.74}\pm 0.00$ $\mathbf{0.96}\pm 0.00$ $\mathbf{0.71}\pm 0.02$ $\mathbf{0.76}\pm 0.01$ $\mathbf{0.95}\pm 0.01$
CytoNet-ViT (1M) P $0.38\pm 0.00$ $0.44\pm 0.00$ $0.79\pm 0.00$ $\mathbf{0.71}\pm 0.00$ $\mathbf{0.75}\pm 0.00$ $\mathbf{0.94}\pm 0.00$
CytoNet-ViT (1M) M $0.54\pm 0.00$ $0.60\pm 0.00$ $0.90\pm 0.00$ $0.65\pm 0.05$ $0.70\pm 0.04$ $0.93\pm 0.02$
transfer brain
scratch - - - $0.10\pm 0.08$ $0.19\pm 0.16$ $0.34\pm 0.26$
scratch-ViT - - - $0.15\pm 0.02$ $0.27\pm 0.03$ $0.49\pm 0.04$
SimCLR (200k) $0.10\pm 0.00$ $0.21\pm 0.01$ $0.39\pm 0.01$ $0.11\pm 0.10$ $0.21\pm 0.17$ $0.38\pm 0.30$
SimCLR (1M) $0.08\pm 0.00$ $0.20\pm 0.01$ $0.35\pm 0.01$ $0.14\pm 0.00$ $0.25\pm 0.02$ $0.47\pm 0.04$
SimCLR-ViT (200k) $0.05\pm 0.00$ $0.16\pm 0.01$ $0.26\pm 0.01$ $0.14\pm 0.04$ $0.24\pm 0.05$ $0.48\pm 0.07$
SimCLR-ViT (1M) $0.05\pm 0.00$ $0.17\pm 0.01$ $0.27\pm 0.01$ $0.13\pm 0.04$ $0.24\pm 0.09$ $0.46\pm 0.11$
SupCon $0.22\pm 0.00$ $0.38\pm 0.00$ $0.67\pm 0.00$ $0.17\pm 0.02$ $0.30\pm 0.02$ $0.55\pm 0.03$
SupCon-ViT $0.25\pm 0.00$ $0.41\pm 0.00$ $0.69\pm 0.00$ $0.15\pm 0.02$ $0.28\pm 0.04$ $0.52\pm 0.06$
CytoNet (200k) $0.32\pm 0.00$ $0.49\pm 0.00$ $0.79\pm 0.00$ $0.18\pm 0.02$ $0.33\pm 0.02$ $0.58\pm 0.05$
CytoNet (1M) $0.35\pm 0.00$ $0.52\pm 0.00$ $0.84\pm 0.00$ $0.17\pm 0.00$ $0.33\pm 0.02$ $0.56\pm 0.02$
CytoNet-ViT (200k) $0.32\pm 0.01$ $0.49\pm 0.01$ $0.82\pm 0.02$ $\mathbf{0.30}\pm 0.01$ $\mathbf{0.46}\pm 0.00$ $\mathbf{0.77}\pm 0.01$
CytoNet-ViT (1M) $\mathbf{0.38}\pm 0.00$ $\mathbf{0.56}\pm 0.00$ $\mathbf{0.88}\pm 0.00$ $0.26\pm 0.03$ $0.43\pm 0.03$ $0.71\pm 0.04$
CytoNet-ViT (1M) P $0.23\pm 0.00$ $0.36\pm 0.00$ $0.71\pm 0.00$ $\mathbf{0.30}\pm 0.01$ $\mathbf{0.48}\pm 0.01$ $\mathbf{0.76}\pm 0.01$
CytoNet-ViT (1M) M $0.37\pm 0.00$ $0.53\pm 0.00$ $0.87\pm 0.00$ $0.20\pm 0.02$ $0.36\pm 0.03$ $0.62\pm 0.03$
unseen brain
scratch - - - $0.12\pm 0.09$ $0.24\pm 0.14$ $0.41\pm 0.23$
scratch-ViT - - - $0.14\pm 0.00$ $0.28\pm 0.00$ $0.47\pm 0.01$
SimCLR (200k) $0.10\pm 0.00$ $0.20\pm 0.00$ $0.36\pm 0.01$ $0.13\pm 0.11$ $0.22\pm 0.18$ $0.39\pm 0.32$
SimCLR (1M) $0.08\pm 0.00$ $0.17\pm 0.01$ $0.32\pm 0.01$ $0.16\pm 0.02$ $0.28\pm 0.02$ $0.51\pm 0.03$
SimCLR-ViT (200k) $0.03\pm 0.00$ $0.12\pm 0.00$ $0.21\pm 0.01$ $0.11\pm 0.01$ $0.21\pm 0.04$ $0.42\pm 0.04$
SimCLR-ViT (1M) $0.04\pm 0.00$ $0.13\pm 0.00$ $0.23\pm 0.01$ $0.12\pm 0.03$ $0.24\pm 0.04$ $0.45\pm 0.07$
SupCon $0.23\pm 0.00$ $0.36\pm 0.01$ $0.62\pm 0.00$ $0.18\pm 0.01$ $0.32\pm 0.01$ $0.54\pm 0.01$
SupCon-ViT $0.24\pm 0.00$ $0.38\pm 0.00$ $0.64\pm 0.01$ $0.14\pm 0.04$ $0.27\pm 0.06$ $0.49\pm 0.09$
CytoNet (200k) $\mathbf{0.28}\pm 0.00$ $\mathbf{0.43}\pm 0.00$ $\mathbf{0.73}\pm 0.01$ $0.16\pm 0.01$ $0.31\pm 0.02$ $0.52\pm 0.02$
CytoNet (1M) $0.25\pm 0.00$ $0.37\pm 0.00$ $0.66\pm 0.01$ $0.17\pm 0.03$ $0.32\pm 0.04$ $0.52\pm 0.05$
CytoNet-ViT (200k) $0.24\pm 0.01$ $0.37\pm 0.01$ $0.67\pm 0.01$ $\mathbf{0.23}\pm 0.01$ $\mathbf{0.37}\pm 0.00$ $\mathbf{0.64}\pm 0.01$
CytoNet-ViT (1M) $0.24\pm 0.00$ $0.36\pm 0.00$ $0.64\pm 0.00$ $0.19\pm 0.02$ $0.35\pm 0.02$ $0.57\pm 0.03$
CytoNet-ViT (1M) P $0.18\pm 0.00$ $0.27\pm 0.00$ $0.55\pm 0.00$ $0.20\pm 0.01$ $0.36\pm 0.01$ $0.59\pm 0.02$
CytoNet-ViT (1M) M $0.23\pm 0.00$ $0.36\pm 0.00$ $0.65\pm 0.00$ $0.16\pm 0.01$ $0.29\pm 0.03$ $0.53\pm 0.03$
Table 5 shows macro-F1 scores, top-1 and top-3 accuracy achieved by different models in brain area classification (Section 3.5) on seen, transfer, and unseen subjects. Performance for finetuning (trainable encoder) and linear probing (frozen encoder) are reported.
CytoNet-ViT (1M) P was trained for area classification on the output of the projection layer rather than the backbone. Using linear probing, its classification performance dropped markedly compared to models using the backbone features, indicating that the projection layer discards or reshapes information essential for brain area classification. After finetuning, the model matched or exceeded the original backbone-based performance, especially on transfer and unseen subjects. These improvements, however, likely reflect the additional model capacity rather than the intrinsic utility of the pretrained projection space, making its specific contribution difficult to isolate.
CytoNet-ViT (1M) M was pretrained using coordinates from the ICBM 152 template (ICBM 2009c Nonlinear asymmetric, [97]) instead of MNI Colin 27. MNI Colin 27 was chosen as default because it provides a single-subject template with well-defined cortical landmarks, whereas ICBM 152 offers a population average with better inter-subject correspondence. Coordinates were non-linearly transformed from MNI Colin 27 to ICBM 152 space using siibra-python [144]. The pretrained model was then linearly probed and finetuned for brain area classification. Pretraining on ICBM 152 yielded classification scores slightly below those obtained with Colin27, but overall performance was comparable, suggesting that the choice of template has limited impact on the learned representations. Notably, transforming MNI Colin 27 coordinates to ICBM MNI 152 rather than natively aligning sections to ICBM MNI 152 may introduce additional errors. The observed differences likely reflect residual registration errors in the range of a few hundred micrometers, which are small compared to the millimeter-scale distances used in the loss but may still introduce local misalignments that affect pretraining. Natively aligning histological sections to ICBM 152 would allow a more direct assessment of template choice, but is challenging due to the limited availability of well-defined landmarks for registration.
Tabelle 6: Scores for brain area classification obtained by linear probing of CytoNet-ViT (1M) backbone with cross-validation across transfer brains. Models were pretrained in different training settings, each considering one of the brains B11, B33, B44, B55, B66, B77, B1010, B1212, and B2020 as transfer brain, and the remaining as seen brains. In all cases, B99 was considered as unseen brain. Performance on seen brains and the unseen is largely independent of the brain used for pretraining. Performance on the respective transfer brain varies slightly, which is likely a result of different subsets of areas that were annotated in each brain.
macro-F1 top-1 acc top-3 acc
seen brains $0.70\pm 0.00$ $0.74\pm 0.00$ $0.96\pm 0.00$
transfer brain $0.37\pm 0.08$ $0.57\pm 0.02$ $0.88\pm 0.01$
unseen brain $0.23\pm 0.00$ $0.36\pm 0.00$ $0.64\pm 0.00$
We evaluated the impact of the selected transfer brain by linear probing of CytoNet-ViT (1M) with training data from different sets of brains (Table 6), keeping the unseen brain B99 fixed. Performance for seen brains and the unseen brain was comparable across all choices of transfer brain. Macro-F1 scores for the transfer brain varied somewhat, reflecting differences in the availability and composition of annotated areas across brains. Since macro-F1 is particularly sensitive to missing labels (e.g., a missing label that is incorrectly predicted only once contributes a zero to the average), such variability is expected and does not affect the overall conclusion: our experiments with transfer brain B77 are representative for the proposed approach.
Tabelle 7: Macro-F1 scores for cortical layer segmentation across different models and training fractions. Mean and standard deviation are reported over 5-fold cross-validation for each model, using either linear probing or finetuning. Models were trained on increasing fractions of the training set, and evaluated on a dedicated test set comprising 184 samples.
linear probe 1% (n=7) 5% (n=36) 10% (n=73) 20% (n=146) 100% (n=732)
SimCLR (200k) $0.49\pm 0.02$ $1\pm 0.01 0.66±0$$6\pm 0.01$ $0.67\pm 0.02$ $0.71\pm 0.01$
SimCLR-ViT (1M) $0.35\pm 0.03$ $0\pm 0.01 0.55±0$$5\pm 0.01$ $0.56\pm 0.01$ $0.59\pm 0.00$
CytoNet (200k) $0.59\pm 0.06$ $0.72\pm 0.01$ $\mathbf{0.74}\pm 0.00$ $5\pm 0.00 0.77±0$$7\pm 0.00$
CytoNet-ViT (1M) $\mathbf{0.63}\pm 0.01$ $\mathbf{0.73}\pm 0.00$ $\mathbf{0.74}\pm 0.00$ $5\pm 0.00 0.77±0$$7\pm 0.00$
finetune
scratch $0.15\pm 0.10$ $0.40\pm 0.22$ $0.65\pm 0.02$ $0.72\pm 0.02$ $\mathbf{0.78}\pm 0.00$
scratch-ViT $0.08\pm 0.05$ $0.52\pm 0.10$ $0.65\pm 0.05$ $0.34\pm 0.41$ $0.20\pm 0.19$
SimCLR (200k) $0.15\pm 0.07$ $0.45\pm 0.05$ $0.53\pm 0.07$ $0.69\pm 0.02$ $0.76\pm 0.03$
SimCLR-ViT (1M) $0.11\pm 0.03$ $0.38\pm 0.06$ $0.50\pm 0.14$ $0.56\pm 0.19$ $0.45\pm 0.26$
CytoNet (200k) $0.23\pm 0.11$ $0.57\pm 0.02$ $0.64\pm 0.05$ $0.73\pm 0.01$ $\mathbf{0.78}\pm 0.02$
CytoNet-ViT (1M) $0.05\pm 0.03$ $0.25\pm 0.18$ $0.52\pm 0.14$ $\mathbf{0.78}\pm 0.02$ $0.21\pm 0.33$
7.5 Visualization of data-driven area classification through clustering
Abbildung 9: Visualization of annotation-based clustering of areas Fp1 and Fp2 in B66 based on CytoNet-ViT (1M) features. The surface mesh of the Colin 27 reference template is shown for reference. A: Points were pre-localized using joint annotations of areas Fp1 and Fp2. B: Points were clustered into two clusters using K-means and were assigned to represent areas Fp1 or Fp2 based on visual inspection. C: Color-coding of the alignment between annotations and cluster assignment.
Figure 9 visualizes the annotation-based clustering of areas Fp1 and Fp2 in B66 based on CytoNet-ViT (1M) features. Clusters were manually assigned to represent areas Fp1 or Fp2 based on visual inspection and compared to the reference annotations of the Julich Brain Atlas. The results show a strong alignment between annotations and cluster assignment, with an accuracy of $94.75\%$.
7.6 Investigation of shortcut learning in SimCLR
Abbildung 10: Retrieval-based analysis of features learned by SimCLR (200k), revealing shortcut learning. The left column shows four randomly sampled reference image patches and their corresponding brain area. For each of the selected patches, the three most similar image patches from the dataset are shown, where the similarity is measured by the cosine similarity between their respective SimCLR (200k) features. Image similarity seems to be largely defined by tissue morphology, while being mostly independent of the brain area, and hence, cytoarchitectonic properties. Annotations point out possible confounding factors for the shown examples, including characteristics tissue morphology in rows 1,2,4, or characteristic blood vessel patterns in rows 2 and 3.
Figure 10 shows a retrieval based analysis for the SimCLR (200k) model for four randomly selected reference image patches.
{refcontext}
[sorting=nyt]
Literatur
  • [1] Katrin Amunts, Claude Lepage, Louis Borgeat, Hartmut Mohlberg, Timo Dickscheid, Marc-Étienne Rousseau, Sebastian Bludau, Pierre-Louis Bazin, Lindsay B. Lewis, Ana-Maria Oros-Peusquens, Nadim J. Shah, Thomas Lippert, Karl Zilles and Alan C. Evans “BigBrain: An Ultrahigh-Resolution 3D Human Brain Model” In Science 340.6139, 2013, pp. 1472–1475 DOI: 10.1126/science.1235381
  • [2] Katrin Amunts and Thomas Lippert “Brain Research Challenges Supercomputing” In Science 374.6571, 2021, pp. 1054–1055 DOI: 10.1126/science.abl8519
  • [3] Katrin Amunts, Aleksandar Malikovic, Hartmut Mohlberg, Thorsten Schormann and Karl Zilles “Brodmann’s Areas 17 and 18 Brought into Stereotaxic Space—Where and How Variable?” In NeuroImage 11.1, 2000, pp. 66–84 DOI: 10.1006/nimg.1999.0516
  • [4] Katrin Amunts, Hartmut Mohlberg, Sebastian Bludau and Karl Zilles “Julich-Brain: A 3D Probabilistic Atlas of the Human Brain’s Cytoarchitecture” In Science 369.6506 American Association for the Advancement of Science, 2020, pp. 988 DOI: 10.1126/science.abb4588
  • [5] Katrin Amunts, Axel Schleicher, Uli Bürgel, Hartmut Mohlberg, Harry B.. Uylings and Karl Zilles “Broca’s Region Revisited: Cytoarchitecture and Intersubject Variability” In Journal of Comparative Neurology 412.2, 1999, pp. 319–341 DOI: 10.1002/(SICI)1096-9861(19990920)412:2¡319::AID-CNE10¿3.0.CO;2-7
  • [6] Katrin Amunts et al. “The Coming Decade of Digital Brain Research: A Vision for Neuroscience at the Intersection of Technology and Computing” In Imaging Neuroscience 2, 2024, pp. imag-2–00137 DOI: 10.1162/imag˙a˙00137
  • [7] Frederico A.. Azevedo, Ludmila R.. Carvalho, Lea T. Grinberg, José Marcelo Farfel, Renata E.. Ferretti, Renata E.. Leite, Wilson Jacob Filho, Roberto Lent and Suzana Herculano-Houzel “Equal Numbers of Neuronal and Nonneuronal Cells Make the Human Brain an Isometrically Scaled-up Primate Brain” In The Journal of Comparative Neurology 513.5, 2009, pp. 532–541 DOI: 10.1002/cne.21974
  • [8] Adrien Bardes, Jean Ponce and Yann LeCun “VICReg: Variance-Invariance-Covariance Regularization for Self-Supervised Learning” In International Conference on Learning Representations (ICLR 2022), 2022 arXiv:2105.04906
  • [9] Herbert Bay, Tinne Tuytelaars and Luc Van Gool “Surf: Speeded up Robust Features” In European Conference on Computer Vision Springer, 2006, pp. 404–417
  • [10] S. Bludau, S.B. Eickhoff, H. Mohlberg, S. Caspers, A.R. Laird, P.T. Fox, A. Schleicher, K. Zilles and K. Amunts “Cytoarchitecture, Probability Maps and Functions of the Human Frontal Pole” In NeuroImage 93, 2014, pp. 260–275 DOI: 10.1016/j.neuroimage.2013.05.052
  • [11] Korbinian Brodmann “Vergleichende Lokalisationslehre der Großhirnrinde” Barth, 1909
  • [12] T. Caliński and J Harabasz “A Dendrite Method for Cluster Analysis” In Communications in Statistics 3.1 Taylor & Francis, 1974, pp. 1–27 DOI: 10.1080/03610927408827101
  • [13] Mathilde Caron, Ishan Misra, Julien Mairal, Priya Goyal, Piotr Bojanowski and Armand Joulin “Unsupervised Learning of Visual Features by Contrasting Cluster Assignments” In Advances in Neural Information Processing Systems 33, 2020, pp. 9912–9924 URL: https://proceedings.neurips.cc/paper/2020/file/70feb62b69f16e0238f741fab228fec2-Paper.pdf
  • [14] Tony Chan and Luminita Vese “An Active Contour Model without Edges” In Scale-Space Theories in Computer Vision, Lecture Notes in Computer Science, 1999, pp. 141–151 DOI: 10.1007/3-540-48236-9˙13
  • [15] Ting Chen, Simon Kornblith, Mohammad Norouzi and Geoffrey Hinton “A Simple Framework for Contrastive Learning of Visual Representations” In International Conference on Machine Learning (ICML 2020), 2020, pp. 1597–1607
  • [16] Xinlei Chen and Kaiming He “Exploring Simple Siamese Representation Learning” In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2021, pp. 15750–15758
  • [17] Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, Parker Schuh, Kensen Shi, Sashank Tsvyashchenko, Joshua Maynez, Abhishek Rao, Parker Barnes, Yi Tay, Noam Shazeer, Vinodkumar Prabhakaran, Emily Reif, Nan Du, Ben Hutchinson, Reiner Pope, James Bradbury, Jacob Austin, Michael Isard, Guy Gur-Ari, Pengcheng Yin, Toju Duke, Anselm Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier Garcia, Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David Dohan, Shivani Agrawal, Mark Omernick, Andrew M. Dai, Thanumalayan Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi Wang, Brennan Saeta, Mark Diaz, Orhan Firat, Michele Catasta, Jason Wei, Kathy Meier-Hellstern, Douglas Eck, Jeff Dean, Slav Petrov and Noah Fiedel “PaLM: Scaling Language Modeling with Pathways” In J. Mach. Learn. Res. 24.1, 2023, pp. 240:11324–240:11436
  • [18] Paolo Cignoni, Marco Callieri, Massimiliano Corsini, Matteo Dellepiane, Fabio Ganovelli and Guido Ranzuglia “MeshLab: An Open-Source Mesh Processing Tool” In Eurographics Italian Chapter Conference The Eurographics Association, 2008
  • [19] Timo Dickscheid, S. Bludau, Casey Paquola, Christian Schiffer, Eric Upschulte and Katrin Amunts “Layer-Specific Distributions of Segmented Cells in Different Cytoarchitectonic Regions of BigBrain Iso Cortex”, 2021 URL: https://search.kg.ebrains.eu/instances/f06a2fd1-a9ca-42a3-b754-adaa025adb10
  • [20] Timo Dickscheid, Xiaoyun Gui, Ahmet Simsek, Christian Schiffer, Jean-Francois Mangin, Yann Leprince, Viktor Jirsa, Jan G. Bjaalie, Trygve B. Leergaard, Sebastian Bludau and Katrin Amunts “Siibra: A Software Tool Suite for Realizing a Multilevel Human Brain Atlas from Complex Data Resources”, 2025, pp. 2025.05.20.655042 DOI: 10.1101/2025.05.20.655042
  • [21] Timo Dickscheid, Sarah Haas, Sebastian Bludau, Philipp Glock, Marcel Huysegoms and Katrin Amunts “Towards 3D Reconstruction of Neuronal Cell Distributions from Histological Human Brain Sections” In Future Trends of HPC in a Disruptive Scenario 34 IOS Press, 2019, pp. 223
  • [22] Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold and Sylvain Gelly “An Image Is Worth 16x16 Words: Transformers for Image Recognition at Scale” In International Conference on Learning Representations (ICLR 2020), 2020 arXiv:2010.11929
  • [23] Martin A. Fischler and Robert C. Bolles “Random Sample Consensus: A Paradigm for Model Fitting with Applications to Image Analysis and Automated Cartography” In Communications of the ACM 24.6 ACM New York, NY, USA, 1981, pp. 381–395
  • [24] Vladimir Fonov, Alan C. Evans, Kelly Botteron, C. Almli, Robert C. McKinstry and D. Collins “Unbiased Average Age-Appropriate Atlases for Pediatric Studies” In NeuroImage 54.1, 2011, pp. 313–327 DOI: 10.1016/j.neuroimage.2010.07.033
  • [25] Robert Geirhos, Jörn-Henrik Jacobsen, Claudio Michaelis, Richard Zemel, Wieland Brendel, Matthias Bethge and Felix A. Wichmann “Shortcut Learning in Deep Neural Networks” In Nature Machine Intelligence 2.11 Nature Publishing Group, 2020, pp. 665–673 DOI: 10.1038/s42256-020-00257-z
  • [26] Stefan Geyer, Anders Ledberg, Axel Schleicher, Shigeo Kinomura, Thorsten Schormann, Uli Bürgel, Torkel Klingberg, Jonas Larsson, Karl Zilles and Per E. Roland “Two Different Areas within the Primary Motor Cortex of Man” In Nature 382.6594 Nature Publishing Group, 1996, pp. 805–807 DOI: 10.1038/382805a0
  • [27] Stefan Geyer, Axel Schleicher and Karl Zilles “Areas 3a, 3b, and 1 of Human Primary Somatosensory Cortex: 1. Microstructural Organization and Interindividual Variability” In NeuroImage 10.1, 1999, pp. 63–83 DOI: 10.1006/nimg.1999.0440
  • [28] Jean-Bastien Grill, Florian Strub, Florent Altché, Corentin Tallec, Pierre H. Richemond, Elena Buchatskaya, Carl Doersch, Bernardo Avila Pires, Zhaohan Daniel Guo, Mohammad Gheshlaghi Azar, Bilal Piot, Koray Kavukcuoglu, Rémi Munos and Michal Valko “Bootstrap Your Own Latent: A New Approach to Self-Supervised Learning” In 34th Conference on Neural Information Processing Systems (NeurIPS 2020), 2020 arXiv: http://arxiv.org/abs/2006.07733
  • [29] H. Haug “Remarks on the Determination and Significance of the Gray Cell Coefficient” In The Journal of Comparative Neurology 104.3, 1956, pp. 473–492 DOI: 10.1002/cne.901040306
  • [30] Kaiming He, Haoqi Fan, Yuxin Wu, Saining Xie and Ross Girshick “Momentum Contrast for Unsupervised Visual Representation Learning” In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2020, pp. 9729–9738 arXiv:1911.05722
  • [31] Kaiming He, Xiangyu Zhang, Shaoqing Ren and Jian Sun “Deep Residual Learning for Image Recognition” In 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2016, pp. 770–778
  • [32] Colin J. Holmes, Rick Hoge, Louis Collins, Roger Woods, Arthur W. Toga and Alan C. Evans “Enhancement of MR Images Using Registration for Signal Averaging” In Journal of Computer Assisted Tomography 22.2, 1998, pp. 324–333 URL: https://journals.lww.com/jcat/Fulltext/1998/03000/Enhancement_of_MR_Images_Using_Registration_for.32.aspx
  • [33] Sergey Ioffe and Christian Szegedy “Batch Normalization: Accelerating Deep Network Training by Reducing Internal Covariate Shift” In International Conference on Machine Learning, 2015, pp. 448–456
  • [34] John Jumper, Richard Evans, Alexander Pritzel, Tim Green, Michael Figurnov, Olaf Ronneberger, Kathryn Tunyasuvunakool, Russ Bates, Augustin Žídek, Anna Potapenko, Alex Bridgland, Clemens Meyer, Simon A.. Kohl, Andrew J. Ballard, Andrew Cowie, Bernardino Romera-Paredes, Stanislav Nikolov, Rishub Jain, Jonas Adler, Trevor Back, Stig Petersen, David Reiman, Ellen Clancy, Michal Zielinski, Martin Steinegger, Michalina Pacholska, Tamas Berghammer, Sebastian Bodenstein, David Silver, Oriol Vinyals, Andrew W. Senior, Koray Kavukcuoglu, Pushmeet Kohli and Demis Hassabis “Highly Accurate Protein Structure Prediction with AlphaFold” In Nature 596.7873 Nature Publishing Group, 2021, pp. 583–589 DOI: 10.1038/s41586-021-03819-2
  • [35] Michael Kazhdan, Matthew Bolitho and Hugues Hoppe “Poisson Surface Reconstruction” In Fourth Eurographics Symposium on Geometry Processing, SGP ’06, 2006, pp. 61–70
  • [36] Prannay Khosla, Piotr Teterwak, Chen Wang, Aaron Sarna, Yonglong Tian, Phillip Isola, Aaron Maschinot, Ce Liu and Dilip Krishnan “Supervised Contrastive Learning” In Advances in Neural Information Processing Systems 33 Curran Associates, Inc., 2020, pp. 18661–18673 arXiv: https://proceedings.neurips.cc/paper/2020/hash/d89a66c7c80a29b1bdbab0f2a1a94af8-Abstract.html
  • [37] Ron Kikinis, Steve D. Pieper and Kirby G. Vosburgh “3D Slicer: A Platform for Subject-Specific Image Analysis, Visualization, and Clinical Support” In Intraoperative Imaging and Image-Guided Therapy New York, NY: Springer, 2014, pp. 277–289 DOI: 10.1007/978-1-4614-7657-3˙19
  • [38] Julius Von Kügelgen, Yash Sharma, Luigi Gresele, Wieland Brendel, Bernhard Schölkopf, Michel Besserve and Francesco Locatello “Self-Supervised Learning with Data Augmentations Provably Isolates Content from Style”, 2021 URL: https://openreview.net/forum?id=4pf_pOo0Dt
  • [39] Yann Leprince, Fabrice Poupon, Thierry Delzescaux, Dominique Hasboun, Cyril Poupon and Denis Rivière “Combined Laplacian-equivolumic Model for Studying Cortical Lamination with Ultra High Field MRI (7 T)” In 2015 IEEE 18th International Symposium on Biomedical Imaging (ISBI 2015), 2015, pp. 580–583 DOI: 10.1109/ISBI.2015.7163940
  • [40] Thomas Lewiner, Hélio Lopes, Antônio Wilson Vieira and Geovan Tavares “Efficient Implementation of Marching Cubes’ Cases with Topological Guarantees” In Journal of Graphics Tools 8.2, 2003, pp. 1–15 DOI: 10.1080/10867651.2003.10487582
  • [41] Lindsay Lewis, Claude Lepage, Marc Fournier, Karl Zilles, Katrin Amunts and Alan Evans “BigBrain: Initial Tissue Classification and Surface Extraction” In Annual Meeting of the Organization for Human Brain Mapping, 2014
  • [42] P. Márquez-Neila, L. Baumela and L. Alvarez “A Morphological Approach to Curvature-Based Evolution of Curves and Surfaces” In IEEE Transactions on Pattern Analysis and Machine Intelligence 36.1, 2014, pp. 2–17 DOI: 10.1109/TPAMI.2013.106
  • [43] Leland McInnes, John Healy, Nathaniel Saul and Lukas Großberger “UMAP: Uniform Manifold Approximation and Projection” In Journal of Open Source Software 3.29, 2018, pp. 861 DOI: 10.21105/joss.00861
  • [44] Björn Merker “Silver Staining of Cell Bodies by Means of Physical Development” In Journal of Neuroscience Methods 9.3, 1983, pp. 235–241 DOI: 10.1016/0165-0270(83)90086-9
  • [45] Jonas Ngnawé, Sabyasachi Sahoo, Yann Pequignot, Frédéric Precioso and Christian Gagné “Detecting Brittle Decisions for Free: Leveraging Margin Consistency in Deep Robust Classifiers”, 2024 DOI: 10.48550/arXiv.2406.18451
  • [46] Aaron Oord, Yazhe Li and Oriol Vinyals “Representation Learning with Contrastive Predictive Coding”, 2019 arXiv: http://arxiv.org/abs/1807.03748
  • [47] OpenAI et al. “GPT-4 Technical Report”, 2024 DOI: 10.48550/arXiv.2303.08774
  • [48] Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, Mahmoud Assran, Nicolas Ballas, Wojciech Galuba, Russell Howes, Po-Yao Huang, Shang-Wen Li, Ishan Misra, Michael Rabbat, Vasu Sharma, Gabriel Synnaeve, Hu Xu, Hervé Jegou, Julien Mairal, Patrick Labatut, Armand Joulin and Piotr Bojanowski “DINOv2: Learning Robust Visual Features without Supervision”, 2024 arXiv: http://arxiv.org/abs/2304.07193
  • [49] Nobuyuki Otsu “A Threshold Selection Method from Gray-Level Histograms” In IEEE Transactions on Systems, Man, and Cybernetics 9.1, 1979, pp. 62–66 DOI: 10.1109/TSMC.1979.4310076
  • [50] Bente Pakkenberg, Dorte Pelvig, Lisbeth Marner, Mads J. Bundgaard, Hans Jørgen G. Gundersen, Jens R. Nyengaard and Lisbeth Regeur “Aging and the Human Neocortex” In Experimental Gerontology 38.1, Proceedings of the 6th International Symposium on the Neurobiology and Neuroendocrinology of Aging, 2003, pp. 95–99 DOI: 10.1016/S0531-5565(02)00151-1
  • [51] Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai and Soumith Chintala “PyTorch: An Imperative Style, High-Performance Deep Learning Library” In 33rd Conference on Neural Information Processing Systems (NeurIPS 2019), 2019 arXiv:1912.01703
  • [52] Jonas Pichat, Juan Eugenio Iglesias, Tarek Yousry, Sébastien Ourselin and Marc Modat “A Survey of Methods for 3D Histology Reconstruction” In Medical Image Analysis 46, 2018, pp. 73–105 DOI: 10.1016/j.media.2018.02.004
  • [53] Stephen M. Pizer, E. Amburn, John D. Austin, Robert Cromartie, Ari Geselowitz, Trey Greer, Bart Haar Romeny, John B. Zimmerman and Karel Zuiderveld “Adaptive Histogram Equalization and Its Variations” In Computer vision, graphics, and image processing 39.3 Elsevier, 1987, pp. 355–368
  • [54] Tobias Pohlen, Alexander Hermans, Markus Mathias and Bastian Leibe “Full-Resolution Residual Networks for Semantic Segmentation in Street Scenes” In IEEE Conference on Computer Vision and Pattern Recognition (CVPR’17), 2017 arXiv:1611.08323
  • [55] Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin and Jack Clark “Learning Transferable Visual Models from Natural Language Supervision” In International Conference on Machine Learning PmLR, 2021, pp. 8748–8763 URL: http://proceedings.mlr.press/v139/radford21a
  • [56] D Rivière, D Geffroy, I Denghien, N Souedet and Y Cointepas “BrainVISA: An Extensible Software Environment for Sharing Multimodal Neuroimaging Data and Processing Tools” In NeuroImage 47, 2009, pp. S163 DOI: 10.1016/S1053-8119(09)71720-3
  • [57] Christian Schiffer, Katrin Amunts, Stefan Harmeling and Timo Dickscheid “Contrastive Representation Learning For Whole Brain Cytoarchitectonic Mapping In Histological Human Brain Sections” In 2021 IEEE 18th International Symposium on Biomedical Imaging (ISBI 2021), 2021, pp. 603–606 DOI: 10.1109/ISBI48211.2021.9433986
  • [58] Christian Schiffer, Stefan Harmeling, Katrin Amunts and Timo Dickscheid “2D Histology Meets 3D Topology: Cytoarchitectonic Brain Mapping with Graph Neural Networks” In Medical Image Computing and Computer Assisted Intervention – MICCAI 2021, 2021, pp. 395–404 DOI: 10.1007/978-3-030-87237-3˙38
  • [59] Axel Schleicher, Katrin Amunts, Stefan Geyer, Patricia Morosan and Karl Zilles “Observer-Independent Method for Microstructural Parcellation of Cerebral Cortex: A Quantitative Approach to Cytoarchitectonics” In NeuroImage 9.1, 1999, pp. 165–177 DOI: 10.1006/nimg.1998.0385
  • [60] Hannah Spitzer, Katrin Amunts, Stefan Harmeling and Timo Dickscheid “Compact Feature Representations for Human Brain Cytoarchitecture Using Self-Supervised Learning” In Medical Imaging with Deep Learning (MIDL 2018), 2018 URL: https://openreview.net/forum?id=HyjJ812of
  • [61] Vitaly Surazhsky and Craig Gotsman “Explicit Surface Remeshing” In Eurographics Symposium on Geometry Processing, 2003
  • [62] Philipp Thörnig “JURECA: Data Centric and Booster Modules Implementing the Modular Supercomputing Architecture at Jülich Supercomputing Centre” In Journal of large-scale research facilities JLSRF 7, 2021, pp. A182–A182 DOI: 10.17815/jlsrf-7-182
  • [63] Yonglong Tian, Chen Sun, Ben Poole, Dilip Krishnan, Cordelia Schmid and Phillip Isola “What Makes for Good Views for Contrastive Learning?” In Advances in Neural Information Processing Systems 33, 2020, pp. 6827–6839 arXiv:2005.10243
  • [64] Eric Upschulte, Stefan Harmeling, Katrin Amunts and Timo Dickscheid “Contour Proposal Networks for Biomedical Instance Segmentation” In Medical Image Analysis, 2022 DOI: 10.1016/j.media.2022.102371
  • [65] Cécile Vogt and Oskar Vogt “Allgemeine Ergebnisse unserer Hirnforschung” JA Barth, 1919
  • [66] Constantin Von Economo “Die Cytoarchitektonik der Hirnrinde des erwachsenen Menschen” Wien: Springer, 1925 URL: http://wwwzb.fz-juelich.de/contentenrichment/inhaltsverzeichnisse/bis2009/Economo,%20Constantin%20von-Die%20Cytoarchitektonik-1925.pdf
  • [67] Konrad Wagstyl, Stéphanie Larocque, Guillem Cucurull, Claude Lepage, Joseph Paul Cohen, Sebastian Bludau, Nicola Palomero-Gallagher, Lindsay B. Lewis, Thomas Funck and Hannah Spitzer “BigBrain 3D Atlas of Cortical Layers: Cortical and Laminar Thickness Gradients Diverge in Sensory and Motor Cortices” In PLOS Biology 18.4 Public Library of Science San Francisco, CA USA, 2020, pp. e3000678
  • [68] Konrad Wagstyl, Stéphanie Larocque, Guillem Cucurull, Claude Lepage, Joseph Paul Cohen, Sebastian Bludau, Nicola Palomero-Gallagher, Lindsay B. Lewis, Thomas Funck, Hannah Spitzer, Timo Dickscheid, Paul C. Fletcher, Adriana Romero, Karl Zilles, Katrin Amunts, Yoshua Bengio and Alan C. Evans “Cortical Intensity Profiles Sampled across BigBrain Isocortex (v1.0)” EBRAINS, 2022 DOI: 10.25493/18ED-DS3
  • [69] Konrad Wagstyl, Claude Lepage, Sebastian Bludau, Karl Zilles, Paul C. Fletcher, Katrin Amunts and Alan C. Evans “Mapping Cortical Laminar Structure in the 3d Bigbrain” In Cerebral Cortex 28.7, 2018, pp. 2551–2562
  • [70] Yang You, Igor Gitman and Boris Ginsburg “Large Batch Training of Convolutional Networks”, 2017 DOI: 10.48550/arXiv.1708.03888
  • [71] Jure Zbontar, Li Jing, Ishan Misra, Yann LeCun and Stephane Deny “Barlow Twins: Self-Supervised Learning via Redundancy Reduction” In Proceedings of the 38th International Conference on Machine Learning PMLR, 2021, pp. 12310–12320 URL: https://proceedings.mlr.press/v139/zbontar21a.html
  • [72] Karl Zilles and Katrin Amunts “Individual Variability Is Not Noise” In Trends in Cognitive Sciences 17.4, 2013, pp. 153–155 DOI: 10.1016/j.tics.2013.02.003
Literatur
  • [73] Korbinian Brodmann “Vergleichende Lokalisationslehre der Großhirnrinde” Barth, 1909
  • [74] Cécile Vogt and Oskar Vogt “Allgemeine Ergebnisse unserer Hirnforschung” JA Barth, 1919
  • [75] Constantin Von Economo “Die Cytoarchitektonik der Hirnrinde des erwachsenen Menschen” Wien: Springer, 1925 URL: http://wwwzb.fz-juelich.de/contentenrichment/inhaltsverzeichnisse/bis2009/Economo,%20Constantin%20von-Die%20Cytoarchitektonik-1925.pdf
  • [76] H. Haug “Remarks on the Determination and Significance of the Gray Cell Coefficient” In The Journal of Comparative Neurology 104.3, 1956, pp. 473–492 DOI: 10.1002/cne.901040306
  • [77] T. Caliński and J Harabasz “A Dendrite Method for Cluster Analysis” In Communications in Statistics 3.1 Taylor & Francis, 1974, pp. 1–27 DOI: 10.1080/03610927408827101
  • [78] Nobuyuki Otsu “A Threshold Selection Method from Gray-Level Histograms” In IEEE Transactions on Systems, Man, and Cybernetics 9.1, 1979, pp. 62–66 DOI: 10.1109/TSMC.1979.4310076
  • [79] Martin A. Fischler and Robert C. Bolles “Random Sample Consensus: A Paradigm for Model Fitting with Applications to Image Analysis and Automated Cartography” In Communications of the ACM 24.6 ACM New York, NY, USA, 1981, pp. 381–395
  • [80] Björn Merker “Silver Staining of Cell Bodies by Means of Physical Development” In Journal of Neuroscience Methods 9.3, 1983, pp. 235–241 DOI: 10.1016/0165-0270(83)90086-9
  • [81] Stephen M. Pizer, E. Amburn, John D. Austin, Robert Cromartie, Ari Geselowitz, Trey Greer, Bart Haar Romeny, John B. Zimmerman and Karel Zuiderveld “Adaptive Histogram Equalization and Its Variations” In Computer vision, graphics, and image processing 39.3 Elsevier, 1987, pp. 355–368
  • [82] Stefan Geyer, Anders Ledberg, Axel Schleicher, Shigeo Kinomura, Thorsten Schormann, Uli Bürgel, Torkel Klingberg, Jonas Larsson, Karl Zilles and Per E. Roland “Two Different Areas within the Primary Motor Cortex of Man” In Nature 382.6594 Nature Publishing Group, 1996, pp. 805–807 DOI: 10.1038/382805a0
  • [83] Colin J. Holmes, Rick Hoge, Louis Collins, Roger Woods, Arthur W. Toga and Alan C. Evans “Enhancement of MR Images Using Registration for Signal Averaging” In Journal of Computer Assisted Tomography 22.2, 1998, pp. 324–333 URL: https://journals.lww.com/jcat/Fulltext/1998/03000/Enhancement_of_MR_Images_Using_Registration_for.32.aspx
  • [84] Katrin Amunts, Axel Schleicher, Uli Bürgel, Hartmut Mohlberg, Harry B.. Uylings and Karl Zilles “Broca’s Region Revisited: Cytoarchitecture and Intersubject Variability” In Journal of Comparative Neurology 412.2, 1999, pp. 319–341 DOI: 10.1002/(SICI)1096-9861(19990920)412:2¡319::AID-CNE10¿3.0.CO;2-7
  • [85] Tony Chan and Luminita Vese “An Active Contour Model without Edges” In Scale-Space Theories in Computer Vision, Lecture Notes in Computer Science, 1999, pp. 141–151 DOI: 10.1007/3-540-48236-9˙13
  • [86] Stefan Geyer, Axel Schleicher and Karl Zilles “Areas 3a, 3b, and 1 of Human Primary Somatosensory Cortex: 1. Microstructural Organization and Interindividual Variability” In NeuroImage 10.1, 1999, pp. 63–83 DOI: 10.1006/nimg.1999.0440
  • [87] Axel Schleicher, Katrin Amunts, Stefan Geyer, Patricia Morosan and Karl Zilles “Observer-Independent Method for Microstructural Parcellation of Cerebral Cortex: A Quantitative Approach to Cytoarchitectonics” In NeuroImage 9.1, 1999, pp. 165–177 DOI: 10.1006/nimg.1998.0385
  • [88] Katrin Amunts, Aleksandar Malikovic, Hartmut Mohlberg, Thorsten Schormann and Karl Zilles “Brodmann’s Areas 17 and 18 Brought into Stereotaxic Space—Where and How Variable?” In NeuroImage 11.1, 2000, pp. 66–84 DOI: 10.1006/nimg.1999.0516
  • [89] Thomas Lewiner, Hélio Lopes, Antônio Wilson Vieira and Geovan Tavares “Efficient Implementation of Marching Cubes’ Cases with Topological Guarantees” In Journal of Graphics Tools 8.2, 2003, pp. 1–15 DOI: 10.1080/10867651.2003.10487582
  • [90] Bente Pakkenberg, Dorte Pelvig, Lisbeth Marner, Mads J. Bundgaard, Hans Jørgen G. Gundersen, Jens R. Nyengaard and Lisbeth Regeur “Aging and the Human Neocortex” In Experimental Gerontology 38.1, Proceedings of the 6th International Symposium on the Neurobiology and Neuroendocrinology of Aging, 2003, pp. 95–99 DOI: 10.1016/S0531-5565(02)00151-1
  • [91] Vitaly Surazhsky and Craig Gotsman “Explicit Surface Remeshing” In Eurographics Symposium on Geometry Processing, 2003
  • [92] Herbert Bay, Tinne Tuytelaars and Luc Van Gool “Surf: Speeded up Robust Features” In European Conference on Computer Vision Springer, 2006, pp. 404–417
  • [93] Michael Kazhdan, Matthew Bolitho and Hugues Hoppe “Poisson Surface Reconstruction” In Fourth Eurographics Symposium on Geometry Processing, SGP ’06, 2006, pp. 61–70
  • [94] Paolo Cignoni, Marco Callieri, Massimiliano Corsini, Matteo Dellepiane, Fabio Ganovelli and Guido Ranzuglia “MeshLab: An Open-Source Mesh Processing Tool” In Eurographics Italian Chapter Conference The Eurographics Association, 2008
  • [95] Frederico A.. Azevedo, Ludmila R.. Carvalho, Lea T. Grinberg, José Marcelo Farfel, Renata E.. Ferretti, Renata E.. Leite, Wilson Jacob Filho, Roberto Lent and Suzana Herculano-Houzel “Equal Numbers of Neuronal and Nonneuronal Cells Make the Human Brain an Isometrically Scaled-up Primate Brain” In The Journal of Comparative Neurology 513.5, 2009, pp. 532–541 DOI: 10.1002/cne.21974
  • [96] D Rivière, D Geffroy, I Denghien, N Souedet and Y Cointepas “BrainVISA: An Extensible Software Environment for Sharing Multimodal Neuroimaging Data and Processing Tools” In NeuroImage 47, 2009, pp. S163 DOI: 10.1016/S1053-8119(09)71720-3
  • [97] Vladimir Fonov, Alan C. Evans, Kelly Botteron, C. Almli, Robert C. McKinstry and D. Collins “Unbiased Average Age-Appropriate Atlases for Pediatric Studies” In NeuroImage 54.1, 2011, pp. 313–327 DOI: 10.1016/j.neuroimage.2010.07.033
  • [98] Katrin Amunts, Claude Lepage, Louis Borgeat, Hartmut Mohlberg, Timo Dickscheid, Marc-Étienne Rousseau, Sebastian Bludau, Pierre-Louis Bazin, Lindsay B. Lewis, Ana-Maria Oros-Peusquens, Nadim J. Shah, Thomas Lippert, Karl Zilles and Alan C. Evans “BigBrain: An Ultrahigh-Resolution 3D Human Brain Model” In Science 340.6139, 2013, pp. 1472–1475 DOI: 10.1126/science.1235381
  • [99] Karl Zilles and Katrin Amunts “Individual Variability Is Not Noise” In Trends in Cognitive Sciences 17.4, 2013, pp. 153–155 DOI: 10.1016/j.tics.2013.02.003
  • [100] S. Bludau, S.B. Eickhoff, H. Mohlberg, S. Caspers, A.R. Laird, P.T. Fox, A. Schleicher, K. Zilles and K. Amunts “Cytoarchitecture, Probability Maps and Functions of the Human Frontal Pole” In NeuroImage 93, 2014, pp. 260–275 DOI: 10.1016/j.neuroimage.2013.05.052
  • [101] Ron Kikinis, Steve D. Pieper and Kirby G. Vosburgh “3D Slicer: A Platform for Subject-Specific Image Analysis, Visualization, and Clinical Support” In Intraoperative Imaging and Image-Guided Therapy New York, NY: Springer, 2014, pp. 277–289 DOI: 10.1007/978-1-4614-7657-3˙19
  • [102] Lindsay Lewis, Claude Lepage, Marc Fournier, Karl Zilles, Katrin Amunts and Alan Evans “BigBrain: Initial Tissue Classification and Surface Extraction” In Annual Meeting of the Organization for Human Brain Mapping, 2014
  • [103] P. Márquez-Neila, L. Baumela and L. Alvarez “A Morphological Approach to Curvature-Based Evolution of Curves and Surfaces” In IEEE Transactions on Pattern Analysis and Machine Intelligence 36.1, 2014, pp. 2–17 DOI: 10.1109/TPAMI.2013.106
  • [104] Sergey Ioffe and Christian Szegedy “Batch Normalization: Accelerating Deep Network Training by Reducing Internal Covariate Shift” In International Conference on Machine Learning, 2015, pp. 448–456
  • [105] Yann Leprince, Fabrice Poupon, Thierry Delzescaux, Dominique Hasboun, Cyril Poupon and Denis Rivière “Combined Laplacian-equivolumic Model for Studying Cortical Lamination with Ultra High Field MRI (7 T)” In 2015 IEEE 18th International Symposium on Biomedical Imaging (ISBI 2015), 2015, pp. 580–583 DOI: 10.1109/ISBI.2015.7163940
  • [106] Kaiming He, Xiangyu Zhang, Shaoqing Ren and Jian Sun “Deep Residual Learning for Image Recognition” In 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2016, pp. 770–778
  • [107] Tobias Pohlen, Alexander Hermans, Markus Mathias and Bastian Leibe “Full-Resolution Residual Networks for Semantic Segmentation in Street Scenes” In IEEE Conference on Computer Vision and Pattern Recognition (CVPR’17), 2017 arXiv:1611.08323
  • [108] Yang You, Igor Gitman and Boris Ginsburg “Large Batch Training of Convolutional Networks”, 2017 DOI: 10.48550/arXiv.1708.03888
  • [109] Leland McInnes, John Healy, Nathaniel Saul and Lukas Großberger “UMAP: Uniform Manifold Approximation and Projection” In Journal of Open Source Software 3.29, 2018, pp. 861 DOI: 10.21105/joss.00861
  • [110] Jonas Pichat, Juan Eugenio Iglesias, Tarek Yousry, Sébastien Ourselin and Marc Modat “A Survey of Methods for 3D Histology Reconstruction” In Medical Image Analysis 46, 2018, pp. 73–105 DOI: 10.1016/j.media.2018.02.004
  • [111] Hannah Spitzer, Katrin Amunts, Stefan Harmeling and Timo Dickscheid “Compact Feature Representations for Human Brain Cytoarchitecture Using Self-Supervised Learning” In Medical Imaging with Deep Learning (MIDL 2018), 2018 URL: https://openreview.net/forum?id=HyjJ812of
  • [112] Konrad Wagstyl, Claude Lepage, Sebastian Bludau, Karl Zilles, Paul C. Fletcher, Katrin Amunts and Alan C. Evans “Mapping Cortical Laminar Structure in the 3d Bigbrain” In Cerebral Cortex 28.7, 2018, pp. 2551–2562
  • [113] Timo Dickscheid, Sarah Haas, Sebastian Bludau, Philipp Glock, Marcel Huysegoms and Katrin Amunts “Towards 3D Reconstruction of Neuronal Cell Distributions from Histological Human Brain Sections” In Future Trends of HPC in a Disruptive Scenario 34 IOS Press, 2019, pp. 223
  • [114] Aaron Oord, Yazhe Li and Oriol Vinyals “Representation Learning with Contrastive Predictive Coding”, 2019 arXiv: http://arxiv.org/abs/1807.03748
  • [115] Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai and Soumith Chintala “PyTorch: An Imperative Style, High-Performance Deep Learning Library” In 33rd Conference on Neural Information Processing Systems (NeurIPS 2019), 2019 arXiv:1912.01703
  • [116] Katrin Amunts, Hartmut Mohlberg, Sebastian Bludau and Karl Zilles “Julich-Brain: A 3D Probabilistic Atlas of the Human Brain’s Cytoarchitecture” In Science 369.6506 American Association for the Advancement of Science, 2020, pp. 988 DOI: 10.1126/science.abb4588
  • [117] Mathilde Caron, Ishan Misra, Julien Mairal, Priya Goyal, Piotr Bojanowski and Armand Joulin “Unsupervised Learning of Visual Features by Contrasting Cluster Assignments” In Advances in Neural Information Processing Systems 33, 2020, pp. 9912–9924 URL: https://proceedings.neurips.cc/paper/2020/file/70feb62b69f16e0238f741fab228fec2-Paper.pdf
  • [118] Ting Chen, Simon Kornblith, Mohammad Norouzi and Geoffrey Hinton “A Simple Framework for Contrastive Learning of Visual Representations” In International Conference on Machine Learning (ICML 2020), 2020, pp. 1597–1607
  • [119] Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold and Sylvain Gelly “An Image Is Worth 16x16 Words: Transformers for Image Recognition at Scale” In International Conference on Learning Representations (ICLR 2020), 2020 arXiv:2010.11929
  • [120] Robert Geirhos, Jörn-Henrik Jacobsen, Claudio Michaelis, Richard Zemel, Wieland Brendel, Matthias Bethge and Felix A. Wichmann “Shortcut Learning in Deep Neural Networks” In Nature Machine Intelligence 2.11 Nature Publishing Group, 2020, pp. 665–673 DOI: 10.1038/s42256-020-00257-z
  • [121] Jean-Bastien Grill, Florian Strub, Florent Altché, Corentin Tallec, Pierre H. Richemond, Elena Buchatskaya, Carl Doersch, Bernardo Avila Pires, Zhaohan Daniel Guo, Mohammad Gheshlaghi Azar, Bilal Piot, Koray Kavukcuoglu, Rémi Munos and Michal Valko “Bootstrap Your Own Latent: A New Approach to Self-Supervised Learning” In 34th Conference on Neural Information Processing Systems (NeurIPS 2020), 2020 arXiv: http://arxiv.org/abs/2006.07733
  • [122] Kaiming He, Haoqi Fan, Yuxin Wu, Saining Xie and Ross Girshick “Momentum Contrast for Unsupervised Visual Representation Learning” In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2020, pp. 9729–9738 arXiv:1911.05722
  • [123] Prannay Khosla, Piotr Teterwak, Chen Wang, Aaron Sarna, Yonglong Tian, Phillip Isola, Aaron Maschinot, Ce Liu and Dilip Krishnan “Supervised Contrastive Learning” In Advances in Neural Information Processing Systems 33 Curran Associates, Inc., 2020, pp. 18661–18673 arXiv: https://proceedings.neurips.cc/paper/2020/hash/d89a66c7c80a29b1bdbab0f2a1a94af8-Abstract.html
  • [124] Yonglong Tian, Chen Sun, Ben Poole, Dilip Krishnan, Cordelia Schmid and Phillip Isola “What Makes for Good Views for Contrastive Learning?” In Advances in Neural Information Processing Systems 33, 2020, pp. 6827–6839 arXiv:2005.10243
  • [125] Konrad Wagstyl, Stéphanie Larocque, Guillem Cucurull, Claude Lepage, Joseph Paul Cohen, Sebastian Bludau, Nicola Palomero-Gallagher, Lindsay B. Lewis, Thomas Funck and Hannah Spitzer “BigBrain 3D Atlas of Cortical Layers: Cortical and Laminar Thickness Gradients Diverge in Sensory and Motor Cortices” In PLOS Biology 18.4 Public Library of Science San Francisco, CA USA, 2020, pp. e3000678
  • [126] Katrin Amunts and Thomas Lippert “Brain Research Challenges Supercomputing” In Science 374.6571, 2021, pp. 1054–1055 DOI: 10.1126/science.abl8519
  • [127] Xinlei Chen and Kaiming He “Exploring Simple Siamese Representation Learning” In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2021, pp. 15750–15758
  • [128] Timo Dickscheid, S. Bludau, Casey Paquola, Christian Schiffer, Eric Upschulte and Katrin Amunts “Layer-Specific Distributions of Segmented Cells in Different Cytoarchitectonic Regions of BigBrain Iso Cortex”, 2021 URL: https://search.kg.ebrains.eu/instances/f06a2fd1-a9ca-42a3-b754-adaa025adb10
  • [129] John Jumper, Richard Evans, Alexander Pritzel, Tim Green, Michael Figurnov, Olaf Ronneberger, Kathryn Tunyasuvunakool, Russ Bates, Augustin Žídek, Anna Potapenko, Alex Bridgland, Clemens Meyer, Simon A.. Kohl, Andrew J. Ballard, Andrew Cowie, Bernardino Romera-Paredes, Stanislav Nikolov, Rishub Jain, Jonas Adler, Trevor Back, Stig Petersen, David Reiman, Ellen Clancy, Michal Zielinski, Martin Steinegger, Michalina Pacholska, Tamas Berghammer, Sebastian Bodenstein, David Silver, Oriol Vinyals, Andrew W. Senior, Koray Kavukcuoglu, Pushmeet Kohli and Demis Hassabis “Highly Accurate Protein Structure Prediction with AlphaFold” In Nature 596.7873 Nature Publishing Group, 2021, pp. 583–589 DOI: 10.1038/s41586-021-03819-2
  • [130] Julius Von Kügelgen, Yash Sharma, Luigi Gresele, Wieland Brendel, Bernhard Schölkopf, Michel Besserve and Francesco Locatello “Self-Supervised Learning with Data Augmentations Provably Isolates Content from Style”, 2021 URL: https://openreview.net/forum?id=4pf_pOo0Dt
  • [131] Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin and Jack Clark “Learning Transferable Visual Models from Natural Language Supervision” In International Conference on Machine Learning PmLR, 2021, pp. 8748–8763 URL: http://proceedings.mlr.press/v139/radford21a
  • [132] Christian Schiffer, Katrin Amunts, Stefan Harmeling and Timo Dickscheid “Contrastive Representation Learning For Whole Brain Cytoarchitectonic Mapping In Histological Human Brain Sections” In 2021 IEEE 18th International Symposium on Biomedical Imaging (ISBI 2021), 2021, pp. 603–606 DOI: 10.1109/ISBI48211.2021.9433986
  • [133] Christian Schiffer, Stefan Harmeling, Katrin Amunts and Timo Dickscheid “2D Histology Meets 3D Topology: Cytoarchitectonic Brain Mapping with Graph Neural Networks” In Medical Image Computing and Computer Assisted Intervention – MICCAI 2021, 2021, pp. 395–404 DOI: 10.1007/978-3-030-87237-3˙38
  • [134] Philipp Thörnig “JURECA: Data Centric and Booster Modules Implementing the Modular Supercomputing Architecture at Jülich Supercomputing Centre” In Journal of large-scale research facilities JLSRF 7, 2021, pp. A182–A182 DOI: 10.17815/jlsrf-7-182
  • [135] Jure Zbontar, Li Jing, Ishan Misra, Yann LeCun and Stephane Deny “Barlow Twins: Self-Supervised Learning via Redundancy Reduction” In Proceedings of the 38th International Conference on Machine Learning PMLR, 2021, pp. 12310–12320 URL: https://proceedings.mlr.press/v139/zbontar21a.html
  • [136] Adrien Bardes, Jean Ponce and Yann LeCun “VICReg: Variance-Invariance-Covariance Regularization for Self-Supervised Learning” In International Conference on Learning Representations (ICLR 2022), 2022 arXiv:2105.04906
  • [137] Eric Upschulte, Stefan Harmeling, Katrin Amunts and Timo Dickscheid “Contour Proposal Networks for Biomedical Instance Segmentation” In Medical Image Analysis, 2022 DOI: 10.1016/j.media.2022.102371
  • [138] Konrad Wagstyl, Stéphanie Larocque, Guillem Cucurull, Claude Lepage, Joseph Paul Cohen, Sebastian Bludau, Nicola Palomero-Gallagher, Lindsay B. Lewis, Thomas Funck, Hannah Spitzer, Timo Dickscheid, Paul C. Fletcher, Adriana Romero, Karl Zilles, Katrin Amunts, Yoshua Bengio and Alan C. Evans “Cortical Intensity Profiles Sampled across BigBrain Isocortex (v1.0)” EBRAINS, 2022 DOI: 10.25493/18ED-DS3
  • [139] Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, Parker Schuh, Kensen Shi, Sashank Tsvyashchenko, Joshua Maynez, Abhishek Rao, Parker Barnes, Yi Tay, Noam Shazeer, Vinodkumar Prabhakaran, Emily Reif, Nan Du, Ben Hutchinson, Reiner Pope, James Bradbury, Jacob Austin, Michael Isard, Guy Gur-Ari, Pengcheng Yin, Toju Duke, Anselm Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier Garcia, Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David Dohan, Shivani Agrawal, Mark Omernick, Andrew M. Dai, Thanumalayan Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi Wang, Brennan Saeta, Mark Diaz, Orhan Firat, Michele Catasta, Jason Wei, Kathy Meier-Hellstern, Douglas Eck, Jeff Dean, Slav Petrov and Noah Fiedel “PaLM: Scaling Language Modeling with Pathways” In J. Mach. Learn. Res. 24.1, 2023, pp. 240:11324–240:11436
  • [140] Katrin Amunts et al. “The Coming Decade of Digital Brain Research: A Vision for Neuroscience at the Intersection of Technology and Computing” In Imaging Neuroscience 2, 2024, pp. imag-2–00137 DOI: 10.1162/imag˙a˙00137
  • [141] Jonas Ngnawé, Sabyasachi Sahoo, Yann Pequignot, Frédéric Precioso and Christian Gagné “Detecting Brittle Decisions for Free: Leveraging Margin Consistency in Deep Robust Classifiers”, 2024 DOI: 10.48550/arXiv.2406.18451
  • [142] OpenAI et al. “GPT-4 Technical Report”, 2024 DOI: 10.48550/arXiv.2303.08774
  • [143] Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, Mahmoud Assran, Nicolas Ballas, Wojciech Galuba, Russell Howes, Po-Yao Huang, Shang-Wen Li, Ishan Misra, Michael Rabbat, Vasu Sharma, Gabriel Synnaeve, Hu Xu, Hervé Jegou, Julien Mairal, Patrick Labatut, Armand Joulin and Piotr Bojanowski “DINOv2: Learning Robust Visual Features without Supervision”, 2024 arXiv: http://arxiv.org/abs/2304.07193
  • [144] Timo Dickscheid, Xiaoyun Gui, Ahmet Simsek, Christian Schiffer, Jean-Francois Mangin, Yann Leprince, Viktor Jirsa, Jan G. Bjaalie, Trygve B. Leergaard, Sebastian Bludau and Katrin Amunts “Siibra: A Software Tool Suite for Realizing a Multilevel Human Brain Atlas from Complex Data Resources”, 2025, pp. 2025.05.20.655042 DOI: 10.1101/2025.05.20.655042